### v2

# Bloch-Beilinson conjectures for Hecke characters and Eisenstein cohomology of Picard surfaces

Abstract.
We consider certain families of Hecke characters $\phi$ over a quadratic imaginary field $F$. According to the Bloch-Beilinson conjectures, the order of vanishing of the $L$-function $L(\phi,s)$ at the central point $s=-1$ should be equal to the dimension of the space of extensions of the Tate motive ${\mathbb{Q}}(1)$ by the motive associated with $\phi$.
In this article, we construct candidates for the corresponding extensions of Hodge structures, assuming that the sign of the functional equation of $L(\phi,s)$ is $-1$. This is accomplished through the cohomology of variations of Hodge structures over Picard modular surfaces associated with $F$ and Harder’s theory of Eisenstein cohomology. Furthermore, we demonstrate that these extensions are naturally realized within certain biextensions. We outline a program to compute the biextension height and utilize it to establish the non-triviality of these extensions.
Key words and phrases:
Eisenstein cohomology, Hecke extension, mixed Hodge structures.
2010 Mathematics Subject Classification:
Primary:11G40;14G35; Secondary:11F75;11F70;14D07
1. Introduction
Let $E$ be an elliptic curve over ${\mathbb{Q}}$ and denote by $r$ the rank of the finitely generated abelian group $E({\mathbb{Q}})$. Suppose that the sign $\epsilon(E)$ of the functional equation of the (completed) $L$-function of $E$ is $-1$. One of the striking consequences of the celebrated Gross-Zagier formula ([14]), in this case, is that if $L^{\prime}(E,1)\neq 0$, then $r\geq 1$. Since by the hypothesis on the sign, $L(E,s)$ vanishes at the central point $s=1$, this agrees with the Birch and Swinnerton-Dyer conjecture, which predicts in particular that $r$ should be precisely equal to the order of vanishing of $L(E,s)$ at $s=1$.
In this article, we are interested in certain conjectural analogues of the latter results, when $L(E,s)$ is replaced by the $L$-function $L(\phi,s)$ of an algebraic Hecke character $\phi$ of a quadratic imaginary field $F$, of odd weight $w$. A natural question arises: what kind of arithmetic information is encoded in the order of vanishing of $L(\phi,s)$ at the central point $s=\frac{w+1}{2}$? The integer $w$ is usually referred to as the motivic weight of the $L$-function.
One way to arrive at a reasonable prediction is by rephrasing the conjecture about elliptic curves in a form that lends itself to generalization. Given a rational point $P$ on an elliptic curve $E$ over ${\mathbb{Q}}$, one can consider the open subvariety $U:=E\setminus\{O,P\}$, where $O$ is the zero point. One looks at the localization exact sequence relating cohomology $H^{\bullet}_{Z}(E)$ of $E$ supported on $Z:=\{P,O\}$, cohomology with compact supports $H^{\bullet}_{c}(U)$ of $U$, and cohomology $H^{\bullet}(E)$. Since the divisor $P-O$ gives rise to a class in $H^{2}_{Z}(E)(1)$ that maps to 0 in $H^{2}(E)(1)$, one gets a morphism $\mathds{1}\rightarrow H^{2}_{Z}(E)(1)$ landing in the image of the connecting homomorphism from $H^{1}_{c}(U)(1)$. By pullback, this yields an exact sequence
$0\rightarrow H^{1}(E)(1)\rightarrow E_{P}\rightarrow\mathds{1}\rightarrow 0$
It turns out that the extension $E_{P}$ is non-trivial if and only if $P$ is of infinite order. This holds both in the sense of extensions of Galois representations (when taking étale cohomology) and of Hodge structures (when taking singular cohomology). Thus, one can reformulate the results of Gross-Zagier as saying that if $\epsilon(E)=-1$ and $L^{\prime}(E,1)\neq 0$, then there exists a non-trivial extension of $\mathds{1}$ by $H^{1}(E)(1)$ of geometric origin.
As explained in Section 3.3, the Bloch-Beilinson conjectures imply a generalization of the previous statement to the case of an algebraic Hecke character $\phi$ of $F$.
Conjecture 1.
Let $\phi$ be an algebraic Hecke character of $F$ of odd weight w. If $\epsilon(\phi)=-1$, then there exists a non-trivial extension
$E_{\phi}\in\mathop{\rm Ext}\nolimits^{1}_{\mathop{\rm MHS}\nolimits}(\mathds{1},H_{\phi}(\frac{w+1}{2}))$
of geometric origin.
Here, $\mathop{\rm MHS}\nolimits$ is a suitable category of mixed Hodge structures with coefficients, to be defined in Section 2.9, and $H_{\phi}$ is the rank 1 pure Hodge structure associated with $\phi$. It arises as a subobject of the Hodge structure on the cohomology of a suitable abelian variety over $F$. The property of being of geometric origin will be made precise in Section 3.3; one can think of it as meaning “arising in the cohomology of an algebraic variety”.
The purpose of this paper is to initiate a program whose goal is confirming the above conjecture, for Hecke characters whose type and whose restriction $\phi_{{\mathbb{Q}}}$ to ${\mathbb{Q}}$ are of a special shape. Denote by $\omega_{F|{\mathbb{Q}}}$ the order two Dirichlet character associated to the extension $F|{\mathbb{Q}}$, and by $|\cdot|_{{\mathbb{I}}_{{\mathbb{Q}}}}$ the idelic norm. We prove the following result.
Theorem 1.
(Corollaries 7.2, 7.3) Let $k$ be a positive integer, and let $\phi$ be a Hecke character of $F$ of type $(k,-k-3)$ or $(-k-3,k)$, satisfying
$\phi_{{\mathbb{Q}}}=\omega_{F|{\mathbb{Q}}}|\cdot|^{3}_{{\mathbb{I}}_{{\mathbb{Q}}}}.$
If $\epsilon(\phi)=-1$, then there exists an extension
$E_{\phi}\in\mathop{\rm Ext}\nolimits^{1}_{\mathop{\rm MHS}\nolimits}(\mathds{1},H_{\phi}(-1))$
of geometric origin.
Note that our hypothesis on $\phi_{{\mathbb{Q}}}$ is equivalent to the very natural assumption that the conjugate inverse of $\phi$ verify
$\bar{\phi}^{-1}=\phi(-3)$
where $(-3)$ is the Tate twist (see Remark 6.1). [^1] We have chosen the formulation in the statement because it is the one more adapted to our approach using Eisenstein cohomology.
Work in progress of the two authors is aimed at showing that, at least when $L^{\prime}(\phi,-1)\neq 0$ (a “Gross-Zagier type” hypothesis), the extension provided by the above theorem is indeed non-trivial. In the rest of this introduction, we will first explain the principle of construction of our extension and the evidence for its non-triviality, then give an outline of the next steps of our program. Finally, we will survey the relation with other related works in the literature.
1.1. Ideas of construction and evidence for non-triviality
The strategy for proving our Theorem 1 is inspired by ideas of Harder ([18]) and makes use of his results in [17]. Starting from the Hecke character $\phi$, one can construct a subspace $\partial H^{2}(\phi)$ in degree 2 boundary cohomology, with coefficients in an appropriate local system, of a Picard modular surface $\mathrm{S}_{K}$, that is, a non-projective Shimura variety, attached to a ${\mathbb{Q}}$-algebraic group $\mathrm{G}$ such that $\mathrm{G}({\mathbb{R}})\simeq\mathop{\rm SU}\nolimits(2,1)$. Note that boundary cohomology of such a variety is computed in terms of the boundary of a compactification, but is independent of the choice of the latter. Then, one would like to lift elements of $\partial H^{2}(\phi)$ to classes in the cohomology of $\mathrm{S}_{K}$ itself. Through Langlands’ theory of Eisenstein operators, one sees that the desired lifts can be provided by classes represented by Eisenstein series, well defined as soon as their constant term converges. For Hecke characters whose infinity type and whose shape are as in our Theorem, this convergence is ensured precisely when $L(\phi,-1)=0$. Hence, in this case, we find that $\partial H^{2}(\phi)$ is a subspace of Eisenstein cohomology, i.e. of the restriction map $r$ from cohomology of $\mathrm{S}_{K}$ to its boundary cohomology.
A natural way to enforce the vanishing $L(\phi,-1)=0$ is to ask that $\epsilon(\phi)=-1$. Now, Rogawski’s study of the automorphic spectrum of unitary groups in three variables ([29, 30, 31]) shows that exactly under this sign hypothesis, a suitable subspace $H^{2}(\phi)_{!}$ can be constructed inside the interior cohomology of Picard surfaces with coefficients in the local system as before, where interior cohomology is defined as the kernel of the restriction map $r$. The subspace $H^{2}(\phi)_{!}$ is related to a $\mathrm{G}({\mathbb{A}}_{f})$-representation induced from $\phi$, and is non-trivial when $\phi$ has the shape prescribed by our Theorem.
To put the above two pieces together, one exploits the short exact sequence linking interior cohomology, cohomology, and Eisenstein cohomology, and gets an extension
(1) $0\rightarrow H^{2}(\phi)_{!}\rightarrow\bar{E}_{0}\rightarrow\partial H^{2}(\phi)\rightarrow 0$
which is in fact an extension of Hodge structures. Classical results allow us to describe the Hodge structure on the piece on the left, while the one on the piece on the right can be computed thanks to the analysis of the degeneration of variations of Hodge structure to the boundary of the Baily-Borel compactification $\mathrm{S}^{*}_{K}$ of $\mathrm{S}_{K}$ of [1]. It turns out that up to a suitable twist, the above extension is nothing but an extension
$0\rightarrow H_{\phi}(-1)\rightarrow E_{0}\rightarrow\mathds{1}^{\oplus\dim\partial H^{2}(\phi)}\rightarrow 0$
Any choice of a class in $\partial H^{2}(\phi)$ produces then an extension $E_{\phi}$ of geometric origin
(2) $0\rightarrow H_{\phi}(-1)\rightarrow E_{\phi}\rightarrow\mathds{1}\rightarrow 0$
as required to prove our Theorem.
Note also that the two pieces in the above extension have been constructed from modules under the Hecke algebra which have repeating eigenvalues. The fact that the hypothesis $\epsilon(\phi)=-1$ implies the existence of a sub-Hecke module of interior cohomology of Picard surfaces, isomorphic at all finite places either to a induced representation from an Hecke character, or to a subquotient of it, is indeed one of the striking consequences of Rogawski’s work that we crucially exploit. We see this feature of our construction as the main reason to hope that, at least for a careful choice of the class in $\partial H^{2}(\phi)$, such an extension may be non split. Note that by contrast, the mixed Hodge structure on the first cohomology of a modular curve is split precisely because interior and Eisenstein cohomology have different eigenvalues under the action of the Hecke algebra.
1.2. Our program
With notation as in (1), let $\Psi$ be a class in $\partial H^{2}(\phi)$, and denote by $E_{\phi}(\Psi)$ the corresponding extension $E_{\phi}$ as in (2). To confirm Conjecture 1, we need to provide a criterion on $\Psi$ for the non-triviality of $E_{\phi}(\Psi)$, and to check that it is verified for a wise choice of $\Psi$.
As explained in Corollary 7.6, there exists a subspace $\partial H^{1}(\theta\phi)$ of degree-1 boundary cohomology of the Picard surface $\mathrm{S}_{K}$, such that choosing classes $\Psi^{\prime}$ in $\partial H^{1}(\theta\phi)$ yields extensions $E_{\phi}(\Psi^{\prime})$ that are dual to the $E_{\phi}(\Psi)$’s in a precise sense. We aim to exploit this fact for establishing and verifying the desired criterion, through the following steps:
  1. (1)
    the support $\Theta$ of the class $\Psi$, identified with a formal linear combination of cusps in the Baily-Borel compactification of the Picard surface $\mathrm{S}_{K}$, provides an algebraic cycle $\mathfrak{E}_{\phi}(\Psi)$ inducing the extension $E_{\phi}(\Psi)$ through an appropriate Abel-Jacobi map; similarly, there exist algebraic cycles $\mathfrak{E}_{\phi}(\Psi^{\prime})$ inducing the extensions $E_{\phi}(\Psi^{\prime})$;
  2. (2)
    there is a well-defined height pairing which can be evaluated on the algebraic cycles $\mathfrak{E}_{\phi}(\Psi)$, $\mathfrak{E}_{\phi}(\Psi^{\prime})$, whose non-vanishing implies that the extension $E_{\phi}(\Psi)$ is non-trivial;
  3. (3)
    the above height pairing is a (finite) sum of local contributions for each place of $F$, and the contribution from the place at infinity is expressed as the biextension height ([15], [9]) of a biextension $B_{\phi}(\Psi,\Psi^{\prime})$ of Hodge structures having $E_{\phi}(\Psi)$ and $E_{\phi}(\Psi)^{\prime}$ as subquotients;
  4. (4)
    the computation of the biextension height of $B_{\phi}(\Psi,\Psi^{\prime})$ can be carried out explicitly and yields information about the non-vanishing of the global height pairing.
This program explains the presence of subsection 7.2 of the present paper, devoted to the construction of the biextensions $B_{\phi}(\Psi,\Psi^{\prime})$. In order to make the computation of their biextension height accessible, one needs to provide explicit representatives of cohomology classes in $\partial H^{2}(\phi)$, by means of differential forms. The construction of such explicit representatives, using the language of Lie algebra cohomology, has recently been achieved in the work [2] by the two authors. The latter should hence be seen as providing one of the crucial ingredients required to carry out the above program. Note that it is precisely in the computation of the biextension height that the term $L^{\prime}(\phi,-1)$ is expected to appear, making its non-vanishing necessary (and sufficient) for the non-vanishing of the pairing itself, and hence for the non-vanishing of our extension class.
1.3. Related work and further future directions
When $\phi$ is an algebraic Hecke character of odd weight $w$, the Bloch-Beilinson conjectures ([8], [5]) are much more precise than their consequence studied in this paper, i.e. than Conjecture 1. Namely, they say that the order of vanishing of $L(\phi,s)$ at the central point $s=\frac{w+1}{2}$ should be equal to the (conjecturally finite) rank of a group of algebraic cycles, and they predict the value of the first non-trivial Taylor coefficient $L^{*}(\phi,\frac{w+1}{2})$ of $L(\phi,s)$ at the central point, up to an algebraic number.
Note that $s=\frac{w+1}{2}$ is a critical point, in the sense of Deligne [12], and hence there exists a canonical period associated to it as in op. cit.. Moreover, Deligne predicted that when non-zero, the value of $L(\phi,s)$ at such a point should be an algebraic multiple of his period. Deligne’s conjecture was later proven for algebraic Hecke characters of $CM$-fields by Blasius [6] and the recent preprint [21] proposes a proof for general number fields.
When the order of vanishing of $L(\phi,s)$ at the central point is greater than zero, the algebraic cycles whose existence is predicted by Bloch and Beilinson should give rise to non-trivial extensions, both as Hodge structures and Galois representations, through appropriate Abel-Jacobi maps. Moreover, the “transcendental part” of the value $L^{*}(\phi,\frac{w+1}{2})$ should be given by the product of Deligne’s period and of the determinant of an height pairing, and related to the extension class of the associated extension of Hodge structures in a precise way. In this paper, we propose candidates for the latter extensions. The program that we have outlined in the previous section of this introduction should moreover make it possible to compute their (conjecturally non-zero) extension class. Our further goal is of course to show that this class, and the height pairing that we plan to employ for its calculation, have indeed the expected relationship with the value $L^{*}(\phi,\frac{w+1}{2})$.
Galois representation-theoretic extensions have been constructed in [3], under the hypothesis of odd order of vanishing of $L(\phi,s)$ at the central point, for essentially the same kind of Hecke characters $\phi$ of $F$ that we consider here (the case of even order of vanishing is studied in [19]). As here, one crucial input in loc. cit. is represented by Rogawski’s results, and the source of the Galois representations considered in these works is again the cohomology of Picard modular surfaces. On the other hand, in order to obtain the desired extensions, the authors use $p$-adic families of automorphic forms, without looking at the geometry of the boundary of the Baily-Borel compactification and at Eisenstein cohomology. Hence, their extensions are not of geometric origin in our sense. It would be interesting to study the Galois-theoretic counterpart of our constructions, via étale cohomology, and to put it in relation to their methods.
1.4. Structure of the article
In Section 2 we gather all the preliminary facts that we need about the structure of the group $\mathrm{G}$, its representation theory, its parabolic subgroups, and about its associated locally symmetric spaces, i.e. the Picard surfaces $\mathrm{S}_{K}$, together with the variations of Hodge structures carried by them.
In Section 3, we set up our notations for algebraic Hecke characters. Moreover, we recall the basic facts about their $L$-functions and their associated motives that we need in order to formulate the Bloch-Beilinson conjecture and to explain how Conjecture 1 follows from them.
In Section 4, we recall the structure of the Baily-Borel and Borel-Serre compactifications of Picard modular surfaces, Moreover, we revisit in the detail the computation of boundary cohomology of Picard surfaces and of its associated Hodge structure, using $\mathrm{S}^{*}_{K}$ ([1]), adding the complements we need. We proceed then by recalling in detail the contents of Harder’s paper [17] on which we rely crucially: in subsection 4.4 we record the description of boundary cohomology, using $\bar{\mathrm{S}}_{K}$, in terms of induced representations from Hecke characters, and in Section 5 we explain the description of Eisenstein cohomology, where the $L$-functions of the aforementioned Hecke characters play an essential role. We conclude the section with the construction of the first piece of our extensions, the space $\partial H^{2}(\phi)$.
In Section 6, we explain how to use Rogawski’s work in order to construct the second piece needed for our extensions, i.e. the space $H^{2}(\phi)_{!}$. The sought-for extensions are then constructed in Section 7, whose final subsection is devoted to showing how our extensions fit into suitable biextensions. These, as explained before, will provide the crucial input for our future work towards the proof of Conjecture 1.
2. Preliminaries
The first part of this section fixes notation about the ground field $F$ and the algebraic group $\mathrm{G}$ that will be in force throughout the whole paper. In the following parts, we discuss the root system and Weyl group of $\mathrm{G}$, we fix a parametrization of its irreducible representations, and we explain the structure of its parabolic subgroups. Finally, we introduce the main geometric objects of interest: the locally symmetric spaces attached to $\mathrm{G}$, their structure of Shimura varieties, and the variation of Hodge structure associated to representations of $\mathrm{G}$.
2.1. Basic notations and conventions
We fix a quadratic imaginary extension $F$ of ${\mathbb{Q}}$. The only non-trivial element of $\mathop{\rm Gal}\nolimits(F|{\mathbb{Q}})$ will be denoted by $\alpha$. For $x\in F^{\times}$, its norm $x\cdot\alpha(x)$ will be denoted by $|x|^{2}$.
The symbols ${\mathbb{A}}_{{\mathbb{K}}}$ and ${\mathbb{I}}_{{\mathbb{K}}}$ will denote respectively adèles and idèles over a number field ${\mathbb{K}}$ (no subscript if ${\mathbb{K}}={\mathbb{Q}}$). The quadratic character of ${\mathbb{I}}_{F}$ associated by class field theory to the extension $F|{\mathbb{Q}}$ will be denoted by $\omega_{F|{\mathbb{Q}}}$.
We fix as well a 3-dimensional $F$-vector space $V$, and a $F$-valued, non degenerate hermitian form $J$ on $V$, such that $J\otimes_{F}{\mathbb{C}}$ is of signature $(2,1)$. For an extension $\mathbb{K}$ of ${\mathbb{Q}}$, we will denote by $V_{\mathbb{K}}$ the $\mathbb{K}$-vector space $V\otimes_{{\mathbb{Q}}}\mathbb{K}$.
There always exists a $F$-basis of $V$ with respect to which $J$ is represented by a matrix of the form
$\left(\begin{array}[]{ccc}&&1\\ &\beta&\\ 1&&\end{array}\right)$
where $\beta$ is a strictly positive rational number. We call such a basis a parabolic basis. On the other hand, we call standard basis a basis of $V_{\mathbb{{\mathbb{R}}}}$ in which $J\otimes_{F}{\mathbb{C}}$ is represented by the matrix
$\left(\begin{array}[]{ccc}1&&\\ &1&\\ &&-1\end{array}\right)$
A change-of-basis matrix between two such bases is given by
(3) $\gamma:=\left(\begin{array}[]{ccc}\frac{1}{\sqrt{2}}&&\frac{1}{\sqrt{2}}\\ &\sqrt{\beta}&\\ \frac{1}{\sqrt{2}}&&-\frac{1}{\sqrt{2}}\end{array}\right)$
To the datum of $(V,J)$, we attach the algebraic group over ${\mathbb{Q}}$
$\mathrm{G}:=\mathop{\rm GU}\nolimits(V,J)$
characterized by the property that for every ${\mathbb{Q}}$-algebra $R$,
$\mathrm{G}(R)=\{g\in\mathop{\rm GL}\nolimits_{F\otimes_{{\mathbb{Q}}}R}(V\otimes_{{\mathbb{Q}}}R)|\ \exists\ \nu(g)\in R^{\times},J(g\cdot,g\cdot)=\nu(g)J(\cdot,\cdot)\}$
The resulting character $\nu:\mathrm{G}\rightarrow{\mathbb{G}}_{m}$ is called the multiplier.
We will also need to work with the kernel of the multiplier map
$\mathrm{G}^{\nu=1}:=\mathop{\rm U}\nolimits(V,J)$
and with the derived subgroup of $\mathrm{G}$,
$\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}:=\mathop{\rm SU}\nolimits(V,J)$
2.2. Structure Theory of $\mathrm{G}$
In order to describe the structure of $\mathrm{G}$, recall that $V_{F}$ decomposes as
(4) $V_{F}\simeq V^{+}\oplus V^{-}$
where
$V^{+}:=\{v\in V\otimes_{{\mathbb{Q}}}F\ |\ (x\otimes 1)v=(1\otimes x)v\ \forall x\in F\}\,$
$V^{-}:=\{v\in V\otimes_{{\mathbb{Q}}}F\ |\ (x\otimes 1)v=(1\otimes\alpha(x))v\ \forall x\in F\}.$
The action of $\alpha$ on $V_{F}$, $v\otimes f\mapsto v\otimes\alpha(f)$, exchanges $V^{+}$ and $V^{-}$. The restrictions to $V$ of the projections on $V^{+}$ and on $V^{-}$ induce a $F$-linear and a $F$-antilinear isomorphism of ${\mathbb{Q}}$-vector spaces
(5) $V\simeq V^{+},\quad V\simeq V^{-}\,.$
Now fix any $F$-basis of $V$ and denote by the same symbol $J$ the matrix representing $J$. Under the above identifications, by definition of $\mathrm{G}$, the relation
(6) $g_{|V^{-}}=\nu(g)J^{-1}{}^{t}g^{-1}_{|V^{+}}J$
holds for any $g\in\mathrm{G}(F)$. This yields an isomorphism
(7) $\mathrm{G}_{F}\simeq\mathop{\rm GL}\nolimits(V_{F})\times{\mathbb{G}}_{m,F}$
by sending $g\in\mathrm{G}(F)$ to $(g_{|V^{+}},\nu(g))$. It restricts to an isomorphism
(8) $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}_{F}\simeq\mathop{\rm SL}\nolimits(V_{F})$
Now fix a parabolic basis of $V$ (see conventions). Consider the resulting embedding
$\mathrm{G}\hookrightarrow\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}\mathop{\rm GL}\nolimits_{3,F}$
and call ${\mathcal{B}}$ and $\mathcal{T}$ the standard (upper triangular) Borel and maximal torus of $\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}\mathop{\rm GL}\nolimits_{3,F}$. Then,
$\mathrm{B}:={\mathcal{B}}\cap\mathrm{G}$
is a Borel of $\mathrm{G}$, and
$\mathrm{T}:=\mathcal{T}\cap\mathrm{G}$
is a maximal torus, that we call standard. It is isogenous to $\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}{\mathbb{G}}_{m,F}\times\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}{\mathbb{G}}_{m,F}$. Under the isomorphism
$\mathrm{G}_{F}\simeq\mathop{\rm GL}\nolimits_{3,F}\times{\mathbb{G}}_{m,F}$
deduced from (7), we get an identification
(9) $\mathrm{T}_{F}(F)\simeq\left\{(\left(\begin{array}[]{ccc}a&&\\ &a^{-1}b&\\ &&b^{-1}q\end{array}\right),f)|a,b,q,f\in F^{\times}\right\}$
A maximal torus of $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}$ is then given by $\mathrm{T}\cap\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}$. We denote it by
$\mathrm{T}^{\prime}:=\mathrm{T}\cap\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}$
It is isomorphic to $\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}{\mathbb{G}}_{m,F}$, and we fix the isomorphism to be the one described on ${\mathbb{Q}}$-points by [^2] This is the same convention as in [17, p. 573].
(10) $\mathrm{T}^{\prime}({\mathbb{Q}})\simeq\left\{\left(\begin{array}[]{ccc}\alpha(a)&&\\ &\alpha(a)^{-1}a&\\ &&a^{-1}\end{array}\right)|a\in F^{\times}\right\}$
Inside $\mathrm{T}^{\prime}$, a maximal ${\mathbb{Q}}$-split torus $\mathrm{T}^{s}$ of $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}$, isomorphic to ${\mathbb{G}}_{m}$, is then described as
(11) $\mathrm{T}^{s}({\mathbb{Q}})\simeq\left\{\left(\begin{array}[]{ccc}t&&\\ &1&\\ &&t^{-1}\end{array}\right)\ |\ t\in{\mathbb{Q}}^{\times}\right\}$
2.3. Root System
We parametrize characters on the standard maximal torus $\mathrm{T}_{F}$ of $\mathrm{G}_{F}$ by vectors $(k_{1},k_{2},c,r)$ of integers such that
$k_{1}\geq k_{2}\geq 0,\ c=k_{1}+k_{2}\pmod{2},\ r=\frac{c+k_{1}+k_{2}}{2}\pmod{2}$
The parametrization is defined by associating $(k_{1},k_{2},c,r)$ with the character
(12) $\left(\left(\begin{array}[]{ccc}a&&\\ &a^{-1}b&\\ &&b^{-1}q\end{array}\right),f\right)\mapsto a^{k_{1}-k_{2}}b^{k_{2}}q^{\frac{c-(k_{1}+k_{2})}{2}}\cdot f^{-\frac{1}{2}(r+\frac{3c-(k_{1}+k_{2})}{2})}$
Remark 2.1.
The reason for this choice of parametrization comes from Hodge theory, see Remark 2.5.
There is a isomorphism between the subgroup of characters $\lambda$ of $\mathrm{T}_{F}$ of the form
(13) $(k_{1},k_{2},k_{1}+k_{2},(k_{1}+k_{2}))$
and the group of characters of the maximal torus $\mathrm{T}_{F}^{\prime}$ of $\mathrm{G}_{F}^{\mathop{\rm\tiny{der}}\nolimits}$. The isomorphism is provided by the natural factorization of characters of the above shape through $\mathrm{T}^{\prime}_{F}$. We still call the resulting character $\lambda$ and we denote it by $(k_{1},k_{2})$.
The root system $\Phi$ of the split, semisimple group $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}_{F}$ is of type $A_{2}$. Under the previous identifications, the system $\Delta$ of simple roots is the set $\{(1,-1),(1,2)\}$ and the set $\Phi^{+}$ of positive roots is the set $\{(1,-1),(1,2),(2,1)\}$.
The half-sum of positive roots (and its lift to $\mathrm{T}_{F}$ defined via (13)) will be denoted by
(14) $\delta:=\frac{1}{2}\sum_{\alpha\in\Phi^{+}}\alpha=(2,1)$
in what follows.
2.4. Irreducible Representations
Irreducible representations of $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}_{F}$ are identified with irreducible representations of $\mathop{\rm SL}\nolimits_{3,F}$, and are therefore parametrized by characters $(k_{1},k_{2})$ which are dominant, i.e. such that $k_{1}\geq k_{2}\geq 0$. Irreducible representations of $\mathrm{G}_{F}$ are then parametrized by characters $(k_{1},k_{2},c,r)$ such that $(k_{1},k_{2})$ is dominant; the integers $c$ and $r$ then account for a suitable twist by a power of the determinant and a power of the multiplier. We denote by $V_{\lambda}$ the irreducible representation of $\mathrm{G}_{F}$ of highest weight $\lambda$.
The $k$-th symmetric power $\mathop{\rm Sym}\nolimits^{k}V$ of the standard representation $V$ of $\mathrm{G}_{F}^{\mathop{\rm\tiny{der}}\nolimits}$ corresponds to $\lambda=(k,0)$. We denote by $V_{k}$ its lift [^3] See Rmk. 2.6 for an explanation of this choice of lift. to the irreducible representation $V_{(k,0,k,k)}$ of $\mathrm{G}_{F}$. We denote by $V^{\vee}_{k}$ its dual.
Remark 2.2.
The hermitian form $J$ provides a $F$-antilinear isomorphism
$V_{k}\simeq V^{\vee}_{k}$
From this, and from relation (6), one sees that $V^{\vee}_{k}$ corresponds to the irreducible representation $V_{(k,k,0,k)}$.
We will also need to parametrize representations of $\mathrm{T}^{\prime}$. Consider the isomorphism $\mathrm{T}^{\prime}\simeq\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}{\mathbb{G}}_{m,F}$ given by (10) and the natural embedding $\mathrm{T}^{\prime}\rightarrow\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}\mathrm{T}^{\prime}_{F}$. Fix a character $\chi:\mathrm{T}^{\prime}_{F}\rightarrow{\mathbb{G}}_{m,F}$, restricting under the previous embedding to the map given on ${\mathbb{Q}}$-points by
$\left(\begin{array}[]{ccc}\alpha(a)&&\\ &\alpha(a)^{-1}a&\\ &&a^{-1}\end{array}\right)\mapsto a^{\mu}\alpha(a)^{\nu}\,,$
Then we denote by
(15) $I_{\mu,\nu}$
the one-dimensional $F$-vector space on which $F^{\times}$ acts via multiplication by $\chi$.
2.5. Standard $\mathbb{Q}$-Parabolic Subgroups
The only ${\mathbb{Q}}$-conjugacy class of parabolic subgroups of $\mathrm{G}$ is the class of the Borel subgroup $\mathrm{B}$. Indeed, on one hand, it is clear that $\mathrm{B}$ is a ${\mathbb{Q}}$-parabolic. On the other hand, any other parabolic has to stabilize a subspace $W$ inside $V$, hence its orthogonal $W^{\perp}$, and by dimension considerations we see that $W\cap W^{\perp}$ has to be a line. Hence any ${\mathbb{Q}}$-parabolic $\mathrm{P}$ is the stabilizer of an isotropic line, and since there is a transitive action of $\mathrm{G}({\mathbb{Q}})$ on the set of isotropic lines, $\mathrm{P}$ has to be conjugated to $\mathrm{B}$.
The Levi component of the maximal parabolic $\mathrm{B}$ is precisely the torus $\mathrm{T}$. We denote by $\mathrm{U}$ the unipotent radical of $\mathrm{B}$, so that we have an isomorphism
$\mathrm{B}\simeq\mathrm{T}\rtimes\mathrm{U}$
2.6. Kostant Representatives
We will need to use a classical result of Kostant. To state it, fix a split reductive group $\mathcal{G}$ over a field of characteristic zero, with root system $\mathfrak{r}$ and Weyl group $W$. Denote by $\mathfrak{r}^{+}$ the set of positive roots and fix moreover a parabolic subgroup $\mathcal{P}$ with its unipotent radical $\mathcal{U}$. Let $\mathfrak{u}$ be the Lie algebra of $\mathcal{U}$ and $\mathfrak{r}_{\mathcal{U}}$ the set of roots appearing inside $\mathfrak{u}$ (necessarily positive). For every $w\in W$, we define:
$\mathfrak{r}^{+}(w):=\{\alpha\in\mathfrak{r}^{+}|w^{-1}\alpha\notin\mathfrak{r}^{+}\},$
$l(w):=|\mathfrak{r}^{+}(w)|,$
$W^{\prime}:=\{w\in W|\mathfrak{r}^{+}(w)\subset\mathfrak{r}_{\mathcal{U}}\}.$
Then, Kostant’s theorem reads as follows:
Theorem 2.3.
[34, Thm. 3.2.3] Let $\mathcal{V}_{\lambda}$ be the irreducible $\mathcal{G}$-representation of highest weight $\lambda$, and let $\delta$ be the half-sum of the positive roots of $\mathcal{G}$. Then, as $(\mathcal{P}/\mathcal{U})$-representations,
$H^{q}(\mathcal{U},\mathcal{V}_{\lambda})\simeq\bigoplus\limits_{w\in W^{\prime}|l(w)=q}\mathcal{V}^{\mathcal{P}/\mathcal{U}}_{w.(\lambda+\delta)-\delta},$
where $\mathcal{V}^{\mathcal{P}/\mathcal{U}}_{\mu}$ denotes the irreducible $(\mathcal{P}/\mathcal{U})$-representation of highest weight $\mu$.
Coming back to our situation, the Weyl group $W$ of $\mathrm{G}_{F}$ is isomorphic to the symmetric group on 3 elements. Fix a character $\lambda=(k_{1},k_{2},c,r)$ of $\mathrm{T}_{F}$. Then, for an element $w\in W$, one computes that the “twisted” Weyl action
(16) $w\star\lambda:=w.(\lambda+\delta)-\delta$
appearing in Kostant’s theorem is given as follows:
$id\star{\lambda}$ $=\lambda$
(1 $(1\ 2)\star\lambda$ $=(k_{2}-1,k_{1}+1,c,r)$
(2 $(2\ 3)\star\lambda$ $=(k_{1}-k_{2}-1,-k_{2}-2,c-k_{2}-1,r)$
(1 2 $(1\ 2\ 3)\star\lambda$ $=(-k_{2}-3,k_{1}-k_{2},c-k_{2}-1,r)$
(1 3 $(1\ 3\ 2)\star\lambda$ $=(k_{2}-k_{1}-3,-k_{1}-3,c-k_{1}-2,r)$
(1 $(1\ 3)\star\lambda$ $=(-k_{1}-4,k_{2}-k_{1}-2,c-k_{1}-2,r)$
The only element of length 0 is the identity, the elements (1 2)12(1\ 2)( 1 2 ) and (2 3)23(2\ 3)( 2 3 ) have length 1, the elements (1 2 3)123(1\ 2\ 3)( 1 2 3 ) and (1 3 2)132(1\ 3\ 2)( 1 3 2 ) have length 2, and the element (1 3)13(1\ 3)( 1 3 ) has length 3.
Notice that the restricted Weyl group of $\mathrm{G}$ is isomorphic to $\mathbb{Z}_{2}$. We fix a generator of it, whose lift $\theta$ to $\mathrm{G}$ is given by (in a parabolic basis)
(17) $\theta:=\left(\begin{array}[]{ccc}&&-1\\ &-1&\\ -1&&\end{array}\right)$
2.7. The associated symmetric spaces and Shimura varieties
In order to describe the locally symmetric spaces associated with $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}$, observe that $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}({\mathbb{R}})\simeq\mathop{\rm SU}\nolimits(2,1)$. Fix a standard basis of $V_{{\mathbb{R}}}$, consider the resulting embedding $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}({\mathbb{R}})\hookrightarrow\mathop{\rm GL}\nolimits_{3}({\mathbb{C}})$ and write an element $g\in\mathrm{G}({\mathbb{R}})$ in block-matrix form as
$\left(\begin{array}[]{cc}A&B\\ C&D\end{array}\right)$
where $A$ is a 2-by-2 matrix, $B$ and $C$ are respectively a 2-by-1 and a 1-by-2 matrix, and $D$ is a scalar. There is a standard maximal compact subgroup $K_{\infty}$ of $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}({\mathbb{R}})$, given by the subgroup of elements
(18) $\left\{\left(\begin{array}[]{cc}A&\\ &\det(A)^{-1}\end{array}\right)\ |\ A\in\mathrm{U}_{2}({\mathbb{R}})\right\}$
where $\det A\in\mathrm{U}_{1}({\mathbb{R}})\simeq S^{1}$. Then, $K_{\infty}\simeq\mathrm{U}_{2}({\mathbb{R}})$, and each other maximal compact subgroup of $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}({\mathbb{R}})$ is conjugated to $K_{\infty}$.
The symmetric space associated to $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}$ is then
$\mathrm{S}:=\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}({\mathbb{R}})/K_{\infty}$
The Lie group $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}({\mathbb{R}})$ acts transitively by generalized fractional transformations on the “matrix” space
$\mathcal{D}_{2,1}:=\{U\in M_{2,1}({\mathbb{C}})|\bar{U}U-1<0\}$
which is nothing but a complex 2-ball in ${\mathbb{C}}^{2}$. The stabilizer of 0 being precisely $K_{\infty}$, the action defines a diffeomorphism between $\mathrm{S}$ and $\mathcal{D}_{2,1}$. This yields a complex analytic structure on $\mathrm{S}$. Then, to any arithmetic subgroup $\Gamma\subset\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}({\mathbb{Q}})$, one associates the locally symmetric space
$\mathrm{S}_{\Gamma}:=\Gamma\backslash\mathrm{S}$
and the complex analytic structure on $\mathrm{S}$ gives $\mathrm{S}_{\Gamma}$ the structure of a complex analytic variety.
When $\Gamma$ is a congruence subgroup, the above locally symmetric spaces can be seen as (analytifications of connected components of) Shimura varieties. To see this, one fixes a parabolic basis of $V$, considers the Deligne torus ${\mathbb{S}}:=\mathop{\rm Res}\nolimits_{{\mathbb{C}}|{\mathbb{R}}}{\mathbb{G}}_{m,{\mathbb{R}}}$ and defines the morphism $h:{\mathbb{S}}\rightarrow\mathrm{G}_{{\mathbb{R}}}$ on ${\mathbb{R}}$-points by
(19) $x+iy\mapsto\left(\begin{array}[]{ccc}x&0&iy\\ 0&x+iy&0\\ iy&0&x\end{array}\right)$
It defines a pure Shimura datum[24, 3.2] [^4] To make the comparison with Lan’s setting, one has to notice that $i\cdot\big{(}\begin{smallmatrix}&&1\\ &1&\\ i&&\end{smallmatrix}\big{)}\cdot J_{\beta}\cdot\big{(}\begin{smallmatrix}&&-i\\ &1&\\ 1&&\end{smallmatrix}\big{)}=\big{(}\begin{smallmatrix}&&1\\ &i\beta&\\ -1&&\end{smallmatrix}\big{)}$, where the latter matrix represents a skew-hermitian form, and that conjugating the above Shimura datum by $\big{(}\begin{smallmatrix}&&-i\\ &1&\\ 1&&\end{smallmatrix}\big{)}$ gives the complex conjugate of Lan’s Shimura datum (loc. cit., Eq. (3.2.5.4)).). Then, the $\mathrm{G}({\mathbb{R}})$-conjugacy class $X$ of $h$ is in canonical bijection with a disjoint union of copies of $\mathrm{S}$ and therefore acquires a canonical topology. For any compact open subgroup $K\subset\mathrm{G}({\mathbb{A}}_{f})$ which is neat in the sense of [27, Sec. 0.6], the set of double classes
$\mathrm{G}({\mathbb{Q}})\backslash(X\times\mathrm{G}({\mathbb{A}}_{f})/K)$
equipped with its natural topology, is canonically homeomorphic to a finite disjoint union of locally symmetric spaces
(20) $\bigsqcup_{i}\mathrm{S}_{\Gamma_{i}}$
where the $\Gamma_{i}$’s are neat (in particular torsion-free) congruence subgroups of $\mathrm{G}({\mathbb{Q}})$, given by
$\Gamma_{i}=g_{i}Kg_{i}^{-1}\cap\mathrm{G}({\mathbb{Q}})$
for suitable elements $g_{i}\in\mathrm{G}({\mathbb{A}}_{f})$. As a consequence, the above set of double classes acquires a canonical structure of complex analytic space, and can be shown to be isomorphic to the analytification of a canonical smooth and quasi-projective algebraic surface $\mathrm{S}_{K}$, called a Picard surface of level $K$, defined over $F$. It satisfies
$\mathrm{S}_{K}({\mathbb{C}})=\mathrm{G}({\mathbb{Q}})\backslash(X\times\mathrm{G}({\mathbb{A}}_{f})/K)$
Remark 2.4.
Each surface $\mathrm{S}_{K}$ can be described as a moduli space of abelian threefolds equipped with additional structures depending on $K$, amongst which complex multiplication by the ring of integers of $F$. In particular, there exists an universal abelian threefold over $\mathrm{S}_{K}$.
The locally symmetric spaces $\Gamma_{i}\backslash\mathrm{S}$ are identified to the analytifications of smooth, geometrically connected quasi-projective surfaces, defined over an abelian extension of $F$ depending on $K$.
2.8. Local systems on $\mathrm{S}_{K}({\mathbb{C}})$
Fix one of the locally symmetric spaces $S_{\Gamma_{i}}$ arising as in (20), and call it $\mathrm{S}_{\Gamma}$. Its fundamental group is identified to $\Gamma$. Hence, any finite dimensional representation of $\mathrm{G}$, by restriction to $\Gamma$, naturally defines a local system on $\mathrm{S}_{\Gamma}$. For any irreducible representation $V_{\lambda}$ of $\mathrm{G}$ of highest weight $\lambda$, we denote by the same symbol the resulting local system. One gets an isomorphism
(21) $H^{\bullet}(\Gamma,V_{\lambda})\cong H^{\bullet}(\mathrm{S}_{\Gamma},V_{\lambda}),$
The above local systems provide local systems $V_{\lambda}$ on $\mathrm{S}_{K}({\mathbb{C}})$ in the obvious way.
2.9. Hodge structures with coefficients and variations of Hodge structures on $\mathrm{S}_{K}({\mathbb{C}})$
We will denote by $\mathop{\rm MHS}\nolimits$ the category of mixed ${\mathbb{Q}}$-Hodge structures. Moreover, we will need to work with a generalization of this category: for ${\mathbb{K}}$ a finite extension of ${\mathbb{Q}}$, the category $\mathop{\rm MHS}\nolimits_{{\mathbb{K}}}$ of mixed Hodge structures with coefficients in ${\mathbb{K}}$, or of mixed ${\mathbb{K}}$-Hodge structures, is the pseudo-abelian completion of the category whose objects are the same as $\mathop{\rm MHS}\nolimits$ and where the morphisms between objects $A,B$ are defined by $\mathop{\rm Hom}\nolimits_{\mathop{\rm MHS}\nolimits}(A,B)\otimes{\mathbb{K}}$. When we do not want to specify the number field ${\mathbb{K}}$, we will speak of Hodge structures with $\overline{{\mathbb{Q}}}$-coefficients. We leave to the reader the task of guessing and working out the precise definition of the category $\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}}$. There are obvious notions of weights and types for a $\overline{{\mathbb{Q}}}$-Hodge structure. Notice that the types of a pure $\overline{{\mathbb{Q}}}$-Hodge structure do not necessarily satisfy Hodge symmetry.
For any irreducible representation $V_{\lambda}$ of $\mathrm{G}$ of highest weight $\lambda$, the morphism $h$ of (19), defining the pure Shimura datum on $\mathrm{G}$, induces a pure $F$-Hodge structure on $V_{\lambda}$. Conjugate $h$ via $g\in\mathrm{G}({\mathbb{R}})$ to obtain a morphism $h^{g}$ whose image lands in the real points $\mathrm{T}({\mathbb{R}})$ of the standard maximal torus of $\mathrm{G}$. Then, by definition, the weight of the Hodge structure on $V_{\lambda}$ is the unique integer $w$ such that for any $t\in{\mathbb{R}}$,
(22) $\lambda\circ h^{g}(t)=t^{-w}$
Remark 2.5.
A computation shows that if $\lambda=(k_{1},k_{2},c,r)$, then
(23) $w=r$
This property is the reason for our choice of parametrizing characters as in (12). In particular, $V_{k}$ is endowed with a pure $F$-Hodge structure of weight $k$.
By the above considerations and by the theory of Shimura varieties, the local systems $V_{\lambda}$ of the previous paragraph are endowed with the structure of a variation of $F$-Hodge structure.
Remark 2.6.
Remember that $V_{k}$ is a choice of lift to $\mathrm{G}$ of the $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}$-representation $\mathop{\rm Sym}\nolimits^{k}V$. Our choice is the one making the variation of Hodge structure $V_{k}$ isomorphic to the $k$-symmetric power of the relative $H^{1}$ of the universal abelian threefold over $\mathrm{S}_{K}$ (Remark 2.4).
By general mixed Hodge theory, the cohomology groups $H^{\bullet}(\mathrm{S}_{K},V_{\lambda})$ are endowed with a canonical mixed $F$-Hodge structure.
2.10. Cohomological dimension
We can estimate the vanishing of the cohomology groups considered in the previous subsection. In fact, following [10], we know that, since $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}$ is a semisimple group over ${\mathbb{Q}}$, for any torsion-free arithmetic subgroup $\Gamma$ of $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}({\mathbb{Q}})$ its cohomological dimension satisfies
$\mathrm{cd}\Gamma=\dim\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}({\mathbb{R}})-\dim K_{\infty}-\mathrm{rank}_{\mathbb{Q}}\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits},$
In our case, we have $\dim\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}({\mathbb{R}})=8$, $\dim K_{\infty}=\dim\mathrm{U}(2)=4$ and $\mathrm{rank}_{\mathbb{Q}}\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}=1$, so that $\mathrm{cd}\Gamma=3$. Therefore, for any irreducible representation $V_{\lambda}$ of $\mathrm{G}$, we get $H^{q}(\mathrm{S}_{\Gamma},V_{\lambda})=0$ for all $q\geq 4$.
3. Algebraic Hecke characters and their motives
In this section, we define algebraic Hecke characters, survey the properties of their $L$-functions, and recall basic facts about the associated motives. We then discuss the Bloch-Beilinson conjectures for Hecke characters, and explain how they imply the existence of non-trivial extensions of the unit Hodge structure by a twist of the Hodge structure associated to an algebraic Hecke character.
3.1. Algebraic Hecke characters
An algebraic Hecke character on a ${\mathbb{Q}}$-torus $\mathcal{T}$ is a continuous homomorphism
$\mathcal{T}({\mathbb{Q}})\backslash\mathcal{T}({\mathbb{A}})\rightarrow{\mathbb{C}}^{\times}\,.$
We will often refer to a Hecke character on $\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}{\mathbb{G}}_{m,F}$ as a Hecke character of $F$.
A Hecke character on $\mathcal{T}$ is algebraic if there exist a character
$\chi:\mathcal{T}\rightarrow{\mathbb{G}}_{m}$
such that the component at infinity $\phi_{\infty}$ of $\phi$ is equal to the character on $\mathcal{T}({\mathbb{C}})$ induced by $\chi^{-1}$. In this case, we say that $\phi$ is of type $\chi$. When $\mathcal{T}=\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}{\mathbb{G}}_{m,F}$, $\phi$ being algebraic amounts to saying that there exist integers $(a,b)$ such that $\phi_{\infty}$ coincides with
××$\begin{array}[]{ccc}{\mathbb{C}}^{\times}&\rightarrow&{\mathbb{C}}^{\times}\\ z&\mapsto&z^{-a}\bar{z}^{-b}\end{array}$
In this case, we say that $(a,b)$ is the infinity type of $\phi$, and $w(\phi):=a+b$ is its weight. Note that the finite part $\phi_{f}$ of an algebraic Hecke character takes values in $\overline{{\mathbb{Q}}}^{\times}$, see e.g. [16, Sec. 2.5]. More precisely, it takes values in a number field $F^{\prime}$, called its field of values.
In our case, algebraic Hecke characters on $\mathrm{T}^{\prime}$ will be identified to Hecke characters of $F$ through the isomorphism (10). As at the end of subsection 2.4, consider the natural embedding $\mathrm{T}^{\prime}\rightarrow\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}\mathrm{T}^{\prime}_{F}$. Fix a character $\chi:\mathrm{T}^{\prime}_{F}\rightarrow{\mathbb{G}}_{m,F}$, and write it $\chi=(\kappa_{1},\kappa_{2})$ in the parametrization of subsection 2.3. Then $\chi$ restricts, under the previous embedding, to the map given on ${\mathbb{R}}$-points by
$\left(\begin{array}[]{ccc}\bar{z}&&\\ &\bar{z}^{-1}z&\\ &&z^{-1}\end{array}\right)\mapsto z^{\kappa_{2}}\bar{z}^{\kappa_{1}-\kappa_{2}}\,,$
With these conventions, the weight of an Hecke character $\phi$ of $F$ of type $\chi=(\kappa_{1},\kappa_{2})$ is
$w(\phi)=\kappa_{1}$
and its infinity type is $(\kappa_{2},\kappa_{1}-\kappa_{2})$.
For later use, let us recall some basic properties of the $L$-function $L(\phi,s)$ of an algebraic Hecke character $\phi$ of $F$. The region in which $L(\phi,s)$ is expressed as an absolutely convergent Euler product is the half-plane $\mbox{Re}\ s>1+\frac{w(\phi)}{2}$. In this region, the $L$-function is therefore holomorphic, and moreover known to be non-zero. Multiplying the Euler product by appropriate $\Gamma$-factors, we obtain a completed L-function $\Lambda(\phi,s)$ which admits a meromorphic continuation to the whole of ${\mathbb{C}}$. The function $\Lambda(\phi,s)$ and the meromorphically continued $L(\phi,s)$ are entire unless $\phi(\cdot)=|\cdot|_{{\mathbb{I}}_{F}}^{t}$ for some $t\in{\mathbb{C}}$; in the latter case, $\Lambda(\phi,s)$ has first-order poles precisely in $s=-t$ and $s=1-t$, and $L(\phi,s)$ has a unique first-order pole in $s=1-t$. Finally, there exists a holomorphic function $\epsilon(s)$ such that $\Lambda(\phi,s)$ satisfies the functional equation
(24) $\Lambda(\phi,s)=\epsilon(s)\Lambda(\phi^{-1},1-s)\,.$
To an algebraic Hecke character of $F$ with field of values $F^{\prime}$, one associates a pure Hodge structure $H_{\phi}$ with coefficients in $F^{\prime}$, and a compatible system of $\ell$-adic Galois representations $(M_{\phi,\ell})_{\ell}$, for $\ell$ varying over the primes. The Hodge structure $H_{\phi}$ is of rank 1 over $F^{\prime}$. Its weight is equal to $w(\phi)$, and its Hodge type is the infinity type of $\phi$.
3.2. Motives for Hecke characters
By [13, Prop. 3.5], there exists a Chow motive $M_{\phi}$ with coefficients in a finite extension of $F^{\prime}$, whose realizations are the Hodge structure $H_{\phi}$ and the compatible system $(M_{\phi,\ell})_{\ell}$. One then has $L(\phi,s)=L(M_{\phi},s)$. The absolute Hodge motive corresponding to $M_{\phi}$ is the one attached to $\phi$ in [32, Ch.1, Thm. 4.1]. Recall that by [13], for every abelian variety $A$, there exists a canonical Chow motive $h^{i}(A)$ realising to the $i$-th cohomology of $A$ in any Weil cohomology theory. Then, the motives $M_{\phi}$ have the following form: if the weight $w$ of $\phi$ is positive, there exist an abelian variety $A_{\phi}$ over $F$, with complex multiplication by a commutative semisimple algebra $T$ over $F$, and an idempotent element $e_{\phi}$ of $T^{\otimes w}$, such that
$M_{\phi}=(h^{1}(A_{\phi})^{\otimes w},e_{\phi})=(h^{w}(A^{w}_{\phi}),e_{\phi})$
Moreover, there exists an isomorphism ([22, Thm. 3.3.1])
$h^{w}(A_{\phi}^{w})\simeq\mathop{\rm Sym}\nolimits^{w}h^{1}(A_{\phi}^{w})$
Since
$\mathop{\rm Sym}\nolimits^{w}h^{1}(A_{\phi}^{w})\simeq\mathop{\rm Sym}\nolimits^{w}h^{1}(A_{\phi})^{\oplus w}$
one gets a projector $e_{1,w}$ defining $h^{1}(A_{\phi})^{\otimes w}$ as a subobject of $h^{w}(A_{\phi}^{w})$. Hence, one can rewrite
$M_{\phi}=(h^{w}(A_{\phi}^{w}),e_{\phi}e_{1,w})$
For any embedding $F\hookrightarrow{\mathbb{C}}$, we can consider the singular cohomology spaces $H^{*}(A^{w}_{\phi})$, with coefficients in a suitable number field, of the analytification of $A^{w}_{\phi}$ with respect to the chosen embedding. This space has a canonical pure Hodge structure of weight $w$, with coefficients in the same number field. The Hodge structure $H_{\phi}$ can be then described as
(25) $H_{\phi}=e_{\phi}e_{1,w}H^{w}(A_{\phi}^{w})$
If the weight of $\phi$ is negative, then the above descriptions hold for the duals of $M_{\phi}$ and $H_{\phi}$.
3.3. Bloch-Beilinson conjectures for Hecke characters
Let $\phi$ be an algebraic Hecke character of $F$, of positive odd weight $w$. When $L(\phi,s)$ vanishes at the central point $s=\frac{w+1}{2}$, the Bloch-Beilinson conjectures predict the value, up to an algebraic number, of the first non-trivial Taylor coefficient of $L(\phi,s)$ at $s=\frac{w+1}{2}$. Let us formulate their prediction precisely (following [26] and [33]).
Consider the motive $M_{\phi}=(h^{1}(A_{\phi})^{\otimes w},e_{\phi})$ attached to $\phi$ as in the previous subsection. Recall that since the integer $\frac{w+1}{2}$ is critical for $M_{\phi}$, Deligne has used the comparison isomorphism between the Betti and de Rham realizations of the weight $-1$ motive $M_{\phi}(\frac{w+1}{2})$ to attach to it a canonical non-zero period $c^{+}(M_{\phi})\in{\mathbb{R}}$. He conjectured that if the value $L(M_{\phi},\frac{w+1}{2})$ does not vanish, then it should be equal to an algebraic multiple of $c^{+}(M_{\phi})$. He offered no statement when $L(\phi,s)$ has a zero at the central point. In order to express what is believed to happen in the latter case, consider the Chow group $CH^{n}(X)$ of a smooth, projective variety $X$ and its subgroup $CH^{n}(X)_{0}$ formed by cycles homologically equivalent to 0, for $n\in\{0,\dots,\dim X\}$. Bloch and Beilinson conjectured the existence of a canonical, non-degenerate height pairing
$h^{n}_{X}:CH^{n}(X)_{0}\otimes{\mathbb{Q}}\times CH^{\dim X+1-n}(X)_{0}\otimes{\mathbb{Q}}\rightarrow{\mathbb{R}}$
Using it, they formulated a conjecture, whose statement we want to express in the special case of $L(\phi,s)$.
In order to do so, we have to extend the conjecture to the Chow groups of our motive $M_{\phi}$. Recall that for $X,n$ as above, there is a natural action of $CH^{\dim X}(X\times X)$ on $CH^{n}(X)$, preserving $CH^{n}(X)_{0}$. Denote by $\Gamma_{*}$ the action of an element $\Gamma\in CH^{\dim X}(X\times X)$. For such a $\Gamma$, the pairing $h^{n}_{X}$ restricts to a pairing, also conjecturally non-degenerate
$h^{n}_{X,\Gamma}:\Gamma_{*}CH^{n}(X)_{0}\otimes{\mathbb{Q}}\times\Gamma_{*}CH^{\dim X+1-n}(X)_{0}\otimes{\mathbb{Q}}\rightarrow{\mathbb{R}}$
Then, writing $L^{*}(\phi,\frac{w+1}{2})$ for the first non-trivial coefficient of $L(\phi,s)$ at $s=\frac{w+1}{2}$, the conjecture reads as follows.
Conjecture 3.1.
Let $M_{\phi}=(h^{w}(A^{w}_{\phi}),e_{\phi}e_{1,w})$ be the Chow motive attached to an algebraic Hecke character $\phi$ of $F$ of positive odd weight $w$. Then
  1. (1)
    $\mathop{\rm\tiny{ord}}\nolimits_{s=\frac{w+1}{2}}L(\phi,s)=\dim_{{\mathbb{Q}}}(e_{\phi}e_{1,w})_{*}CH^{\frac{w+1}{2}}(A^{w}_{\phi})_{0}\otimes{\mathbb{Q}}$
  2. (2)
    Denoting by $\sim_{\overline{{\mathbb{Q}}}}$ the relation of coinciding up to multiplication by a non-zero algebraic number,
    $L^{*}(\phi,\frac{w+1}{2})\sim_{\overline{{\mathbb{Q}}}}c^{+}(M_{\phi})\det(h^{\frac{w+1}{2}}_{A_{\phi}^{w},e_{\phi}e_{1,w}})$
The conjecture extends in the obvious way to the case of negative weight $w$.
Now consider the Abel-Jacobi map
$e_{\phi}e_{1,w}CH^{\frac{w+1}{2}}(A^{w}_{\phi})_{0}\rightarrow\mathop{\rm Ext}\nolimits^{1}_{\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}}}(\mathds{1},e_{\phi}e_{1,w}H^{w}(A^{w}_{\phi})(\frac{w+1}{2}))$
Since $A_{\phi}^{w}$ is defined over a number field, this map is conjectured to be injective ([5, 5.6]). Using the description (25) of $H_{\phi}$, it takes the form
(26) $e_{\phi}e_{1,w}CH^{\frac{w+1}{2}}(A^{w}_{\phi})_{0}\rightarrow\mathop{\rm Ext}\nolimits^{1}_{\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}}}(\mathds{1},H_{\phi}(\frac{w+1}{2}))$
Let $\epsilon(\phi)$ be the sign of the functional equation of $\phi$. The conjectural injectivity of the map (26), point (1) of Conjecture 3.2, and the fact that $\epsilon(\phi)=-1$ implies that $L(\phi,s)$ vanishes at the central point, lead to the following (considerably weaker and less precise) conjecture.
Conjecture 3.2.
Let $\phi$ be an algebraic Hecke character of $F$ of odd weight w. If $\epsilon(\phi)=-1$, then there exists a non-trivial extension
$E_{\phi}\in\mathop{\rm Ext}\nolimits^{1}_{\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}}}(\mathds{1},H_{\phi}(\frac{w+1}{2}))$
of geometric origin.
A way of making the above property of being of geometric origin precise is the following. There is a triangulated, $\overline{{\mathbb{Q}}}$-linear tensor category $\mathop{\rm DM}\nolimits_{c}(F)_{\overline{{\mathbb{Q}}}}$ of constructible motives over $F$, containing the category of Chow motives as a full subcategory. It is equipped with a realization functor
$R_{H}:\mathop{\rm DM}\nolimits_{c}(F)_{\overline{{\mathbb{Q}}}}\rightarrow D^{b}(\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}})$
Then, we say that an object $E$ of $\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}}$ is of geometric origin if it belongs to the full Tannakian subcategory of $\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}}$ generated by the cohomology objects of objects belonging to the image of $R_{H}$.
4. Boundary cohomology
In this section, we fix one of the Picard modular surfaces $\mathrm{S}_{K}$ introduced in Section 2.7, and let $\mathrm{S}_{\Gamma}$ be a connected component of $\mathrm{S}_{K}({\mathbb{C}})$. We aim to describe the boundary cohomology of $\mathrm{S}_{K}({\mathbb{C}})$ with coefficients in local systems coming from representations of the underlying group $\mathrm{G}$, together with their mixed Hodge structures and $\mathrm{G}({\mathbb{A}}_{f})$-actions. To do so, we will consider two different compactifications, the Baily-Borel compactification $\mathrm{S}^{*}_{\Gamma}$ (a complex projective variety) and the Borel-Serre compactification $\overline{\mathrm{S}}_{\Gamma}$ (a real analytic manifold with corners).
4.1. Boundary of the Baily-Borel compactification
For facts recalled in this paragraph, we refer to [28, Sects. 3.6-3.7]. Any ${\mathbb{Q}}$-conjugacy class of parabolics in $\mathrm{G}$ provides a set of strata of the boundary of the Baily-Borel compactification of $\mathrm{S}_{\Gamma}$. More precisely, given a representative $\mathrm{P}$ of such a conjugacy class, there is a canonical reductive subgroup $\mathrm{M}$ of the Levi component of $\mathrm{P}$, and a canonical conjugacy class of morphisms
(27) $h_{\mathrm{M}}:{\mathbb{S}}\rightarrow\mathrm{M}_{{\mathbb{R}}}$
providing a Shimura datum attached to $\mathrm{M}$. Each of the strata of $\mathrm{S}_{\Gamma}^{*}$ corresponding to $\mathrm{P}$ is then a connected component of a Shimura variety associated with this Shimura datum.
In our case, the only ${\mathbb{Q}}$-conjugacy class of parabolics in $\mathrm{G}$ is the one of the Borel $\mathrm{B}$. The Levi component of $\mathrm{B}$ is the torus $\mathrm{T}$, and hence, the positive dimensional, reductive subgroup $\mathrm{M}$ is a torus as well. Thus, the Baily-Borel boundary is a disjoint union of 0-dimensional strata, called cusps.
4.2. Boundary of the Borel-Serre compactification
For facts recalled in this paragraph, we refer to [17, pp. 567-570]. The Borel-Serre compactification $\overline{\mathrm{S}}_{\Gamma}$ of $\mathrm{S}_{\Gamma}$ admits a canonical projection $\pi$ to the Baily-Borel one. Let us describe, in our situation, the fiber of $\pi$ over each cusp $p$ of $\mathrm{S}^{*}_{\Gamma}$. To do so, let us write
$\mathrm{U}^{\prime}:=\mathrm{U}\cap\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}$
Denote by $A_{\mathrm{G}}$ the identity component of the real points of the maximal ${\mathbb{Q}}$-split subtorus $\mathrm{T}^{s}$ of $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}$. From equation (10), we see that there is an isomorphism of Lie groups $\mathrm{T}^{\prime}({\mathbb{R}})\simeq{\mathbb{C}}^{\times}$ which by equation (11) restricts to $A_{\mathrm{G}}\simeq{\mathbb{R}}^{\times}_{>0}.$
Define ${}^{0}T$ as the Lie group $A_{\mathrm{G}}\backslash\mathrm{T}^{\prime}({\mathbb{R}})$. It is isomorphic to $S^{1}$. Denote by $K^{T}_{\infty}$ the intersection with ${}^{0}T$ of a maximal compact subgroup of $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}({\mathbb{R}})$. By conjugating the maximal compact subgroup of equation (18) via the change of basis matrix given in (3), one sees that $K^{T}_{\infty}$ coincides with ${}^{0}T$ itself. On the other hand, denote $\Gamma_{\mathrm{U}}:=\Gamma\cap\mathrm{U}^{\prime}({\mathbb{R}})$. The general theory tells us that the space $\pi^{-1}(p)$ has the structure of a fibration, with base a quotient of ${}^{0}T\backslash K^{T}_{\infty}$ (hence a point in our case), and with fiber diffeomorphic to the nilmanifold
$\mathrm{S}_{\mathrm{U}_{\Gamma}}:=\Gamma_{\mathrm{U}}\backslash\mathrm{U}^{\prime}({\mathbb{R}})$
This means that in our case, $\pi^{-1}(p)\simeq\mathrm{S}_{\mathrm{U}_{\Gamma}}$. Now $\mathrm{U}^{\prime}$ is inserted into an exact sequence
$0\rightarrow\mathrm{W}\rightarrow\mathrm{U}^{\prime}\rightarrow\widetilde{\mathrm{U}^{\prime}}\rightarrow 0$
where $\mathrm{W}$ is the commutator subgroup of $\mathrm{U}^{\prime}$, isomorphic to ${\mathbb{G}}_{a}$, and $\widetilde{\mathrm{U}^{\prime}}$ is isomorphic to $\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}{\mathbb{G}}_{a,F}$. Since we are working with a torsion-free $\Gamma$, this yields a diffeomorphism of $\mathrm{S}_{\mathrm{U}_{\Gamma}}$ with a $S^{1}$-fiber bundle over an elliptic curve.
4.3. Boundary cohomology and its Hodge structure
Let $V_{\lambda}$ be the irreducible representation of $\mathrm{G}$ of highest weight $\lambda$. Let $\widetilde{\mathrm{S}_{\Gamma}}$ be a compactification of $\mathrm{S}_{\Gamma}$ and let $j:\mathrm{S}_{\Gamma}\hookrightarrow\widetilde{\mathrm{S}_{\Gamma}}$ be the corresponding open immersion, with boundary $i:\partial\widetilde{\mathrm{S}_{\Gamma}}\hookrightarrow\widetilde{\mathrm{S}_{\Gamma}}$. Boundary cohomology of $\mathrm{S}_{\Gamma}$ with coefficients in $V_{\lambda}$ is defined as hypercohomology of a complex of sheaves over $\partial\widetilde{\mathrm{S}_{\Gamma}}$:
$\partial H^{n}(\mathrm{S}_{\Gamma},V_{\lambda}):=\mathbb{H}^{n}(\partial\widetilde{\mathrm{S}_{\Gamma}},i^{*}Rj_{*}V_{\lambda})$
By proper base change, it does not depend on the choice of the compactification $\widetilde{\mathrm{S}_{\Gamma}}$. There is a natural long exact sequence
(28) $\cdots\rightarrow H_{c}^{n}(\mathrm{S}_{\Gamma},V_{\lambda})\rightarrow H^{n}(\mathrm{S}_{\Gamma},V_{\lambda})\rightarrow\partial H^{n}(\mathrm{S}_{\Gamma},V_{\lambda})\rightarrow\cdots$
By choosing as compactification the Baily-Borel compactification $S_{\Gamma}^{*}$, mixed Hodge theory endows the spaces $\partial H^{\bullet}(\mathrm{S}_{\Gamma},V_{\lambda})$ with a canonical mixed Hodge structure, and the above long exact sequence becomes one of mixed Hodge structures.
The following Proposition gives a complete description of the mixed Hodge structure on $\partial H^{\bullet}(\mathrm{S}_{\Gamma},V_{\lambda})$, using the spaces $I_{\mu,\nu}$ introduced in (15). It is equivalent to the main result of [1], but we have chosen to write it in detail using our parametrizations and notations and to provide an outline of its proof for the convenience of the reader.
Proposition 4.1.
Let $\lambda=(k_{1},k_{2},c,r)$. Let $m$ be the number of cusps of $\mathrm{S}_{\Gamma}$. Then, for any integer $n$, the mixed Hodge structures $\partial H^{n}(\mathrm{S}_{\Gamma},V_{\lambda})$ are semisimple, and described as follows:
  1. (1)
    the space $\partial H^{n}(\mathrm{S}_{\Gamma},V_{\lambda})$ is trivial for $n$ outside the set $\{0,\dots,3\}$.
  2. (2)
    For $n=0$,
    $\partial H^{0}(\mathrm{S}_{\Gamma},V_{\lambda})\cong I_{k_{2},k_{1}-k_{2}}^{\oplus m}$
[…]
    $\mbox{\emph{weight}:}\ r-k_{1}$
    $\mbox{\emph{type}:}\ \left(-\frac{k_{1}+k_{2}+c-2r}{4},-\frac{3k_{1}-k_{2}-c-2r}{4}\right)$
  3. (3)
    For $n=1$,
    $\partial H^{1}(\mathrm{S}_{\Gamma},V_{\lambda})\simeq I_{k_{1}+1,-k_{1}+k_{2}-2}^{\oplus m}\oplus I_{-k_{2}-2,k_{1}+1}^{\oplus m}$
[…]
    $\mbox{\emph{weight}:}\ (r+1)-k_{2}$
    $\mbox{\emph{type}:}\ \left(-\frac{k_{1}+k_{2}+c-2r}{4},-\frac{-k_{1}+3k_{2}-4-c-2r}{4}\right)$
[…]
    $\mbox{\emph{weight}:}\ (r+1)-(k_{1}-k_{2})$
    $\mbox{\emph{type}:}\ \left(-\frac{k_{1}-3k_{2}-4+c-2r}{4},-\frac{3k_{1}-k_{2}-c-2r}{4}\right)$
  4. (4)
    For $n=2$,
    $\partial H^{2}(\mathrm{S}_{\Gamma},V_{\lambda})\simeq I_{k_{1}-k_{2},-k_{1}-3}^{\oplus m}\oplus I_{-k_{1}-3,k_{2}}^{\oplus m}$
[…]
    $\mbox{\emph{weight}:}\ (r+2)+k_{2}+1$
    $\mbox{\emph{type}:}\ \left(-\frac{k_{1}-3k_{2}-4+c-2r}{4},-\frac{-k_{1}-k_{2}-8-c-2r}{4}\right)$
[…]
    $\mbox{\emph{weight}:}\ (r+2)+(k_{1}-k_{2})+1$
    $\mbox{\emph{type}:}\ \left(-\frac{-3k_{1}+k_{2}-8+c-2r}{4},-\frac{-k_{1}+3k_{2}-4-c-2r}{4}\right)$
  5. (5)
    For $n=3$
    $\partial H^{3}(\mathrm{S}_{\Gamma},V_{\lambda})\simeq I_{-k_{1}-4,-k_{2}-2}^{\oplus m}$
[…]
    $\mbox{\emph{weight}:}\ (r+3)+k_{1}+1$
    $\mbox{\emph{type}:}\ \left(-\frac{-3k_{1}+k_{2}-8+c-2r}{4},-\frac{-k_{1}-k_{2}-8-c-2r}{4}\right)$
Proof.
The assertion of point (1) follows from the vanishing discussed in paragraph 2.10 and from the long exact sequence (28).
To compute boundary cohomology in degree $n\in\{0,\dots,3\}$, let us first observe that since $\partial\mathrm{S}^{*}_{\Gamma}$ is 0-dimensional,
$\partial H^{n}(\mathrm{S}_{\Gamma},V_{\lambda})\simeq i^{*}R^{n}j_{*}V_{\lambda}$
This is also an isomorphism of mixed Hodge structures, once we consider the mixed Hodge structure on $i^{*}R^{n}j_{*}V_{\lambda}$ given by the theory of mixed Hodge modules. Now, for any cusp $c\in\partial\mathrm{S}^{*}_{\Gamma}$, write $i_{c}$ for the composition of $i$ with the closed immersion of $c$ into $\partial\mathrm{S}^{*}_{\Gamma}$. We then get
$\partial H^{n}(\mathrm{S}_{\Gamma},V_{\lambda})\simeq\bigoplus\limits_{c\in\partial\mathrm{S}^{*}_{\Gamma}}i_{c}^{*}R^{n}j_{*}V_{\lambda}$
Now consider the subgroup $\mathrm{M}$ of $\mathrm{T}$, and its associated Shimura datum $h_{M}$, appearing in (27). The general formalism of [27, Ch. 4] has been spelled out in this special case in [1, Lemma 3.6, Lemma 3.8]. The outcome is that the map $h_{M}$ is defined on ${\mathbb{C}}$-points by
(29) $(z_{1},z_{2})\mapsto(\left(\begin{array}[]{ccc}z_{1}z_{2}&0&0\\ 0&z_{1}&0\\ 0&0&1\end{array}\right),z_{1}z_{2})\,$
Through $h_{M}$, any representation of $\mathrm{M}_{{\mathbb{R}}}$ acquires a semisimple Hodge structure.
By [11, Thm. 2.9], there is an isomorphism of mixed Hodge structures
$i_{c}^{*}R^{n}j_{*}V_{\lambda}\simeq\bigoplus\limits_{p+q=n}H^{p}(\overline{H}_{C},H^{q}(\mathrm{U}_{F},V_{\lambda}))$
where $\overline{H}_{C}$ is a certain neat arithmetic subgroup of $\mathrm{T}/\mathrm{M}$, and the Hodge structure on the spaces $H^{p}(\overline{H}_{C},H^{q}(\mathrm{U}_{F},V_{\lambda}))$ is the one obtained through (29) by considering them as $\mathrm{M}$-representations, via restriction of the natural $\mathrm{T}$-representation.
Since $\mathrm{T}$ is isogenous to $\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}{\mathbb{G}}_{m,F}\times\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}{\mathbb{G}}_{m,F}$, any neat arithmetic subgroup of it is trivial, hence the same holds for neat arithmetic subgroups of its quotients. So, the above formula becomes
$i_{c}^{*}R^{n}j_{*}V_{\lambda}\simeq H^{n}(\mathrm{U}_{F},V_{\lambda})$
Notice that in particular, the Hodge structure on $i_{c}^{*}R^{n}j_{*}V_{\lambda}$ is independent of the cusp $c$, so that if $m$ is the number of cusps, then as Hodge structures
$\partial H^{n}(\mathrm{S}_{\Gamma},V_{\lambda})\simeq H^{n}(\mathrm{U}_{F},V_{\lambda})^{\oplus m}$
By applying Kostant’s theorem 2.3, we see that
$H^{n}(\mathrm{U}_{F},V_{\lambda})\simeq\bigoplus\limits_{w\in W^{\prime}|l(w)=n}V^{\mathrm{T}_{F}}_{w.(\lambda+\delta)-\delta},$
where $V^{\mathrm{T}_{F}}_{w.(\lambda+\delta)-\delta}$ is the one-dimensional representation on which the split torus $\mathrm{T}_{F}$ acts via the character $w.(\lambda+\delta)-\delta$. Denoting the latter character by $\chi$, and expressing it in our parametrization by
$\chi=(\kappa_{1},\kappa_{2},\gamma,\rho)$
a computation shows that through $\chi\circ h_{M}$, the element $(z_{1},z_{2})\in{\mathbb{S}}({\mathbb{C}})$ acts on $V^{\mathrm{T}_{F}}_{\chi}$ via multiplication by
(30) $z_{1}^{\frac{\kappa_{1}+\kappa_{2}}{4}+\frac{1}{4}\gamma-\frac{1}{2}\rho}z_{2}^{\frac{3\kappa_{1}-\kappa_{2}}{4}-\frac{1}{4}\gamma-\frac{1}{2}\rho}\,.$
So, the opposites of the exponents of $z_{1}$ and $z_{2}$ appearing above give us the type of the one-dimensional $F$-Hodge structure carried by $V^{\mathrm{T}_{F}}_{\chi}$.
On the other hand, using the isomorphism (10), an element $a\in F^{\times}$ acts on $V^{\mathrm{T}_{F}}_{\chi}$ via multiplication by
$a^{\kappa_{2}}\alpha(a)^{\kappa_{1}-\kappa_{2}}$
so that we get an identification, as $\mathrm{T}^{\prime}$-representations,
$V^{\mathrm{T}_{F}}_{\chi}=I_{\kappa_{2},\kappa_{1}-\kappa_{2}}$
By plugging in the explicit formulae, provided in subsection 2.6, for the different $\chi$’s appearing in Kostant’s theorem, assertions (2)-(5) of the statement are then proven. ∎
Remark 4.2.
  1. (1)
    The assertions in Proposition 4.1 concerning the weights of the Hodge structures of interest can also be found, with our same parametrizations, in [35, proof of Theorem 3.8]. In our way of writing the weights, we have imitated loc. cit. in emphasizing the deviation from purity (i.e., the difference, in a fixed degree $i$, between the actual weight and the pure weight $r+i$). As op. cit. shows, controlling this difference is important for being able to construct motives for cuspidal automorphic forms on $\mathrm{G}$.
  2. (2)
    The topological description of the boundary cohomology spaces $\partial H^{\bullet}(\mathrm{S}_{\Gamma},V_{\lambda})$, with no computation of their Hodge structure (since [11] was not available at that time) was already determined in [17, Eq. (2.1.2)]. In the notation of loc. cit., the elements $w_{\alpha}$, $w_{\bar{\alpha}}$, $w_{\bar{\alpha}}w_{\alpha}$, $w_{\alpha}w_{\bar{\alpha}}$ and $\theta$ correspond respectively to our (2 3)23(2\ 3)( 2 3 ), (1 2)12(1\ 2)( 1 2 ), (1 2 3)123(1\ 2\ 3)( 1 2 3 ), (1 3 2)132(1\ 3\ 2)( 1 3 2 ) and (1 3)13(1\ 3)( 1 3 ).
  3. (3)
    For his analysis of boundary cohomology in [17], Harder chose to work with the Borel-Serre compactification $\overline{\mathrm{S}}_{\Gamma}$. Its crucial property is that the open immersion
    $\mathrm{S}_{\Gamma}\hookrightarrow\overline{\mathrm{S}}_{\Gamma}$
    is an homotopy equivalence, so that the sheaf $V_{\lambda}$ canonically extends to a sheaf on $\overline{\mathrm{S}}_{\Gamma}$, denoted by the same symbol. Hence, there is no need to compute hypercohomology: the spaces $\partial H^{n}(\mathrm{S}_{\Gamma},V_{\lambda})$ are canonically isomorphic to $H^{n}(\partial\overline{\mathrm{S}}_{\Gamma},V_{\lambda})$. As a consequence, for computing boundary cohomology associated with higher-rank groups underlying a Shimura datum, the Borel-Serre compactification has a strong computational advantage over the Baily-Borel one. In our situation, the boundary structure being as simple as possible, determining boundary cohomology as vector spaces with either choice is equally easy. The Baily-Borel compactification gives us a way to describe their canonical mixed Hodge structure. In the following section, we will recall how the Borel-Serre one provides a description of boundary cohomology in terms of automorphic forms.
In the following two corollaries, we specialize Proposition 4.1 to the case of $V_{\lambda}$ being one of the two local systems $V_{k}$ or $V_{k}^{\vee}$ defined in subsection 2.4.
Corollary 4.3.
For $n\in\{0,\dots,3\}$, the Hodge structure on the boundary cohomology spaces $\partial H^{n}(\mathrm{S}_{\Gamma},V_{k})$ is described as follows.
  1. (1)
    For $n=0$,
    $\partial H^{0}(\mathrm{S}_{\Gamma},V_{k})\cong I_{0,k}^{\oplus m}$
[…]
    weight: $\mbox{\emph{weight}:}\ 0$
    $\mbox{\emph{type}:}\ (0,0)$
  2. (2)
    For $n=1$,
    $\partial H^{1}(\mathrm{S}_{\Gamma},V_{k})\simeq I_{k+1,-k-2}^{\oplus m}\oplus I_{-2,k+1}^{\oplus m}$
[…]
    $\mbox{\emph{weight}:}\ k+1$
    $\mbox{\emph{type}:}\ (0,k+1)$
[…]
    weight: $\mbox{\emph{weight}:}\ 1$
    $\mbox{\emph{type}:}\ (1,0)$
  3. (3)
    For $n=2$,
    $\partial H^{2}(\mathrm{S}_{\Gamma},V_{k})\simeq I_{k,-k-3}^{\oplus m}\oplus I_{-k-3,0}^{\oplus m}$
[…]
    $\mbox{\emph{weight}:}\ (k+2)+1$
    $\mbox{\emph{type}:}\ (1,k+2)$
[…]
    $\mbox{\emph{weight}:}\ (k+2)+k+1$
    $\mbox{\emph{type}:}\ (k+2,k+1)$
  4. (4)
    For $n=3$
    $\partial H^{3}(\mathrm{S}_{\Gamma},V_{k})\simeq I_{-k-4,-2}^{\oplus m}$
[…]
    $\mbox{\emph{weight}:}\ (k+3)+k+1$
    $\mbox{\emph{type}:}\ (k+2,k+2)$
Using Rmk. 2.2, one sees:
Corollary 4.4.
For $n\in\{0,\dots,3\}$, the Hodge structure on the boundary cohomology spaces $\partial H^{n}(\mathrm{S}_{\Gamma},V^{\vee}_{k})$ has:
  • as weights, the same as those of the Hodge structure on $\partial H^{n}(\mathrm{S}_{\Gamma},V_{k})$;
  • as types, the conjugates of the ones of the Hodge structure on $\partial H^{n}(\mathrm{S}_{\Gamma},V_{k})$.
4.4. Automorphic structure of boundary cohomology
We end this section by describing the automorphic structure of boundary cohomology. For this purpose, we switch to the adelic setting and work with the spaces
$\mathrm{S}_{K}({\mathbb{C}}):={\mathrm{G}}({\mathbb{Q}})\backslash(X\times{\mathrm{G}}({\mathbb{A}}_{f})/K)$
defined in Remark 2.7. For any irreducible representation $V_{\lambda}$ of $\mathrm{G}$ of highest weight $\lambda$, denote by $V_{\lambda,K}$ the corresponding local system on $\mathrm{S}_{K}({\mathbb{C}})$. Letting $K$ vary among the compact open subgroups of ${\mathrm{G}}({\mathbb{A}}_{f})$, one obtains a projective system $(\mathrm{S}_{K}({\mathbb{C}}),V_{\lambda,K})_{K}$ of spaces and sheaves, the cohomology of whose projective limit $(\mathrm{S}({\mathbb{C}}),V_{\lambda})$ is such that
[…]
and is endowed with a canonical structure of ${\mathrm{G}}({\mathbb{A}}_{f})$-module. There is an analogously defined cohomology with compact support $H^{\bullet}_{c}(\mathrm{S}({\mathbb{C}}),V_{\lambda})$, and one defines interior cohomology
(31) $H_{!}^{\bullet}(\mathrm{S}({\mathbb{C}}),V_{\lambda})$
as the image of the natural morphism $H_{c}^{\bullet}(\mathrm{S}({\mathbb{C}}),V_{\lambda})\rightarrow H^{\bullet}(\mathrm{S}({\mathbb{C}}),V_{\lambda})$. Interior cohomology at a finite level is recover by taking $K$-invariants,
(32) $H_{!}^{\bullet}(\mathrm{S}_{K}({\mathbb{C}}),V_{\lambda})=H_{!}^{\bullet}(\mathrm{S}({\mathbb{C}}),V_{\lambda})^{K}$
Denote by $\partial{\mathrm{S}}_{K}({\mathbb{C}})$ the boundary of the Borel-Serre compactification at any level $K$. From it and from the restriction to $\partial{\mathrm{S}}_{K}({\mathbb{C}})$ of the canonically extended $V_{\lambda}$ we obtain an analogous projective system and, in the projective limit, a ${\mathrm{G}}({\mathbb{A}}_{f})$-module that we denote
$\partial H^{\bullet}(\mathrm{S}({\mathbb{C}}),V_{\lambda})$
Again, boundary cohomology at a finite level is recovered by taking $K$-invariants,
$\partial H^{\bullet}(\mathrm{S}_{K}({\mathbb{C}}),V_{\lambda})=\partial H^{\bullet}(\mathrm{S}({\mathbb{C}}),V_{\lambda})^{K}$
To describe the “automorphic structure of boundary cohomology” means, in a first approximation, to describe the structure of the ${\mathrm{G}}({\mathbb{A}}_{f})$-module $\partial H^{\bullet}(\mathrm{S}({\mathbb{C}}),V_{\lambda})$. This description will be given in terms of Hecke characters.
Definition 4.5.
Let $\phi$ be an algebraic Hecke character of $\mathrm{T}$ and consider its finite component
$\phi_{f}:\mathrm{T}({\mathbb{A}}_{f})\rightarrow\overline{{\mathbb{Q}}}^{\times}$
Let $\mathrm{T}({\mathbb{A}}_{f})$ act on $\overline{{\mathbb{Q}}}$ via multiplication by $\phi_{f}$. We denote by
¯$\overline{{\mathbb{Q}}}_{\phi}$
the resulting $\mathrm{T}({\mathbb{A}}_{f})$-module. We make it a $\mathrm{B}({\mathbb{A}}_{f})$-module by considering the decomposition $\mathrm{B}=\mathrm{T}\mathrm{U}$ and letting $\mathrm{U}({\mathbb{A}}_{f})$ act trivially.
Recall the “twisted” action $w\star\lambda$ of an element $w\in W$ of the Weyl group of $\mathrm{G}$ on a weight $\lambda$, as in subsection 2.6. Then, the theorem we are interested in reads as follows:
Theorem 4.6.
[17, Thm. 1] There exists a canonical isomorphism of ${\mathrm{G}}({\mathbb{A}}_{f})$-modules
(33) $H^{\bullet}(\partial\mathrm{S}({\mathbb{C}}),V_{\lambda,\overline{{\mathbb{Q}}}})\simeq\bigoplus\limits_{w\in W}\bigoplus_{\mbox{\emph{\tiny{type}}}(\phi)=w\star\lambda}\mathop{\rm Ind}\nolimits_{{\mathrm{B}}({\mathbb{A}}_{f})}^{{\mathrm{G}}({\mathbb{A}}_{f})}\overline{{\mathbb{Q}}}_{\phi}$
Let us now specialize the above Theorem to the case of $V_{\lambda}$ being one of the two local systems $V_{k}$ or $V_{k}^{\vee}$ defined in subsection 2.4. If the type of the restriction to $\mathrm{T}^{\prime}$ of a Hecke character $\phi$ of $\mathrm{T}$ is $\chi=(\kappa_{1},\kappa_{2})$, then we denote [^5] The notation is motivated by the fact that in this case, the infinity type of the restriction of $\phi$ to $\mathrm{T}^{\prime}$ is $(\kappa_{2},\kappa_{1}-\kappa_{2})$ (cfr. subsection 3). by
¯$\overline{{\mathbb{Q}}}_{\kappa_{2},\kappa_{1}-\kappa_{2}}$
the $\mathrm{B}({\mathbb{A}}_{f})$-module ¯$\overline{{\mathbb{Q}}}_{\phi}$. The computation of the characters $w\star\lambda$ provided in subsection 2.6 yields then the following.
Corollary 4.7.
Let $\lambda=(k,0,k,k)$.
Then there are canonical isomorphisms of ${\mathrm{G}}({\mathbb{A}}_{f})$-modules
$H^{0}(\partial\mathrm{S}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})\simeq\bigoplus_{\mbox{\emph{\tiny{type}}}(\phi)=\lambda}\mathop{\rm Ind}\nolimits_{{\mathrm{B}}({\mathbb{A}}_{f})}^{{\mathrm{G}}({\mathbb{A}}_{f})}\overline{{\mathbb{Q}}}_{0,k}$
$H^{1}(\partial\mathrm{S}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})\simeq\bigoplus_{\mbox{\emph{\tiny{type}}}(\phi)=(1\ 2)\star\lambda}\mathop{\rm Ind}\nolimits_{{\mathrm{B}}({\mathbb{A}}_{f})}^{{\mathrm{G}}({\mathbb{A}}_{f})}\overline{{\mathbb{Q}}}_{k+1,-k-2}\oplus\bigoplus_{\mbox{\emph{\tiny{type}}}(\phi)=(2\ 3)\star\lambda}\mathop{\rm Ind}\nolimits_{{\mathrm{B}}({\mathbb{A}}_{f})}^{{\mathrm{G}}({\mathbb{A}}_{f})}\overline{{\mathbb{Q}}}_{-2,k+1}$
$H^{2}(\partial\mathrm{S}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})\simeq\bigoplus_{\mbox{\emph{\tiny{type}}}(\phi)=(1\ 2\ 3)\star\lambda}\mathop{\rm Ind}\nolimits_{{\mathrm{B}}({\mathbb{A}}_{f})}^{{\mathrm{G}}({\mathbb{A}}_{f})}\overline{{\mathbb{Q}}}_{k,-k-3}\oplus\bigoplus_{\mbox{\emph{\tiny{type}}}(\phi)=(1\ 3\ 2)\star\lambda}\mathop{\rm Ind}\nolimits_{{\mathrm{B}}({\mathbb{A}}_{f})}^{{\mathrm{G}}({\mathbb{A}}_{f})}\overline{{\mathbb{Q}}}_{-k-3,0}$
$H^{3}(\partial\mathrm{S}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})\simeq\bigoplus_{\mbox{\emph{\tiny{type}}}(\phi)=(1\ 3)\star\lambda}\mathop{\rm Ind}\nolimits_{{\mathrm{B}}({\mathbb{A}}_{f})}^{{\mathrm{G}}({\mathbb{A}}_{f})}\overline{{\mathbb{Q}}}_{-k-4,-2}$
Remark 4.8.
 The relation between the above Corollary and Corollary 4.3 is the following (a similar relation holds indeed for any system of coefficients $V_{\lambda}$).
Fix a neat compact open subgroup $K\subset\mathrm{G}({\mathbb{A}}_{f})$ and a connected component $\mathrm{S}_{\Gamma}$ of $\mathrm{S}_{K}({\mathbb{C}})$. Then $H^{\bullet}(\mathrm{S}_{\Gamma},V_{k,\overline{{\mathbb{Q}}}})$ and $\partial H^{\bullet}(\mathrm{S}_{\Gamma},V_{k,\overline{{\mathbb{Q}}}})$ get identified respectively to a direct summand of $H^{\bullet}(\mathrm{S}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})^{K}$ and $H^{\bullet}(\partial\mathrm{S}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})^{K}$.
Under the resulting inclusion
$\partial H^{\bullet}(\mathrm{S}_{\Gamma},V_{k,\overline{{\mathbb{Q}}}})\hookrightarrow H^{\bullet}(\partial\mathrm{S}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})^{K}$
the spaces $I_{\mu,\nu,\overline{{\mathbb{Q}}}}:=I_{\mu,\nu}\otimes_{F}\overline{{\mathbb{Q}}}$ on the left-hand side become identified to one-dimensional subspaces of the spaces $(\mathop{\rm Ind}\nolimits_{{\mathrm{B}}({\mathbb{A}}_{f})}^{{\mathrm{G}}({\mathbb{A}}_{f})}\overline{{\mathbb{Q}}}_{\mu,\nu})^{K}$ on the right-hand side. Hence, if we write
$m_{\mu,\nu}:=\dim_{\overline{{\mathbb{Q}}}}(\mathop{\rm Ind}\nolimits_{{\mathrm{B}}({\mathbb{A}}_{f})}^{{\mathrm{G}}({\mathbb{A}}_{f})}\overline{{\mathbb{Q}}}_{\mu,\nu})^{K}$
then we get an isomorphism
(34) $I^{\oplus m_{\mu,\nu}}_{\mu,\nu,\overline{{\mathbb{Q}}}}\simeq(\mathop{\rm Ind}\nolimits_{{\mathrm{B}}({\mathbb{A}}_{f})}^{{\mathrm{G}}({\mathbb{A}}_{f})}\overline{{\mathbb{Q}}}_{\mu,\nu})^{K}$
under which the spaces on the right hand-side become endowed with a pure $\overline{{\mathbb{Q}}}$-Hodge structure, direct sum of copies of the pure, 1-dimensional $\overline{{\mathbb{Q}}}$-Hodge structure $I_{\mu,\nu,\overline{{\mathbb{Q}}}}$.
Corollary 4.9.
Let $\lambda=(k,k,0,k)$. Then the $\mathrm{G}({\mathbb{A}}_{f})$-module structure of the boundary cohomology spaces $H^{\bullet}(\partial\mathrm{S}({\mathbb{C}}),V^{\vee}_{k,\overline{{\mathbb{Q}}}})$ is the same as the one on $H^{\bullet}(\partial\mathrm{S}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})$, up to replacing the infinity types of the Hecke characters appearing in the latter with their conjugates.
5. Eisenstein cohomology
In this section, we fix again one of the Picard modular surfaces $\mathrm{S}_{K}$ introduced in Section 2.7 and we describe its Eisenstein cohomology by recalling the results of the article [17] of Harder. This is used at the end of the section in order to construct the first piece of the extensions which are the object of our main theorem, Theorem 1: a sub-Hodge structure $\partial H^{2}(\phi)$ of Eisenstein cohomology, associated to a suitable Hecke character $\phi$.
5.1. Eisenstein cohomology
Following the notations fixed in Sections 2 and 4, we define $Z:=Z({\mathrm{G}})({\mathbb{R}})^{0}$ and $\tilde{K}_{\infty}:=Z\cdot{K}_{\infty}$. Consider the irreducible representation $V_{\lambda}$ of $\mathrm{G}_{F}$ of highest weight $\lambda$ and denote by $\omega$ its central character. We define the space
$C^{\infty}(\mathop{{\rm G}}\nolimits({\mathbb{Q}})\backslash\mathop{{\rm G}}\nolimits({\mathbb{A}}))(\omega^{-1})$
as the space of those ${\mathbb{C}}$-valued, $C^{\infty}$-functions $f$ on $\mathop{{\rm G}}\nolimits({\mathbb{Q}})\backslash\mathop{{\rm G}}\nolimits({\mathbb{A}})$ which satisfy $f(zg)=\omega^{-1}(z)f(g)$ for any $z\in Z(\tilde{\mathrm{G}})({\mathbb{A}})^{0}$, $g\in\mathop{{\rm G}}\nolimits({\mathbb{A}})$. We define the archimedean component
$C^{\infty}(\mathop{{\rm G}}\nolimits({\mathbb{Q}})\backslash\mathop{{\rm G}}\nolimits({\mathbb{R}}))(\omega^{-1})$
and the non-archimedean component
$C^{\infty}(\mathop{{\rm G}}\nolimits({\mathbb{Q}})\backslash\mathop{{\rm G}}\nolimits({\mathbb{A}}_{f}))(\omega^{-1})$
of this space in an analogous way.
We will also need to consider the subspace
$\mathcal{A}(\mathop{{\rm G}}\nolimits({\mathbb{Q}})\backslash\mathop{{\rm G}}\nolimits({\mathbb{R}}))\subset C^{\infty}(\mathop{{\rm G}}\nolimits({\mathbb{Q}})\backslash\mathop{{\rm G}}\nolimits({\mathbb{A}}))$
of automorphic forms on $\mathrm{G}$.
If we denote by ${\mathfrak{g}}$ the complexified Lie algebra of $\mathop{{\rm G}}\nolimits({\mathbb{R}})$, there is a canonical isomorphism
$H^{\bullet}(S({\mathbb{C}}),V_{\lambda,{\mathbb{C}}})\simeq H^{\bullet}({\mathfrak{g}},\tilde{K}_{\infty},C^{\infty}(\mathop{{\rm G}}\nolimits({\mathbb{Q}})\backslash\mathop{{\rm G}}\nolimits({\mathbb{A}}))(\omega^{-1})\otimes V_{\lambda,{\mathbb{C}}})$
which induces an isomorphism of $\mathop{{\rm G}}\nolimits({\mathbb{A}}_{f})$-modules
$H^{\bullet}(S({\mathbb{C}}),V_{\lambda,{\mathbb{C}}})\simeq C^{\infty}(\mathop{{\rm G}}\nolimits({\mathbb{Q}})\backslash\mathop{{\rm G}}\nolimits({\mathbb{A}}_{f}))(\omega^{-1})\otimes H^{\bullet}({\mathfrak{g}},\tilde{K}_{\infty},C^{\infty}(\mathop{{\rm G}}\nolimits({\mathbb{Q}})\backslash\mathop{{\rm G}}\nolimits({\mathbb{R}}))(\omega^{-1})\otimes V_{\lambda,{\mathbb{C}}})\,.$
This section aims to exploit the latter description to determine Eisenstein cohomology, i.e. the image of the map of $\mathop{{\rm G}}\nolimits({\mathbb{A}}_{f})$-modules
(36) $r:H^{\bullet}(\mathrm{S}({\mathbb{C}}),V_{\lambda})\rightarrow\partial H^{\bullet}(\mathrm{S}({\mathbb{C}}),V_{\lambda})\,.$
[…]
(37) $I_{\phi}:=\mathop{\rm Ind}\nolimits_{\mathrm{B}({\mathbb{A}}_{f})}^{\mathrm{G}({\mathbb{A}}_{f})}\overline{{\mathbb{Q}}}_{\phi}$
Moreover, we define the Harish-Chandra module
$I_{\phi,\infty}:=\left\{f:\mathop{{\rm G}}\nolimits({\mathbb{R}})\rightarrow{\mathbb{C}}\ |\ f\ \text{is}\ \tilde{K}_{\infty}-\text{finite and s.t.}f(bg)=\phi_{\infty}(b)f(g)\ \forall b\in\mathrm{B}({\mathbb{R}}),g\in\mathop{{\rm G}}\nolimits({\mathbb{R}})\right\}\,.$
We will then consider the space
$I_{\phi}^{*}:=\mathop{\rm Ind}\nolimits_{\mathrm{B}({\mathbb{A}})}^{\mathrm{G}({\mathbb{A}})}\phi=I_{\phi,{\mathbb{C}}}\otimes I_{\phi,\infty}$
5.2. The Eisenstein operator and its constant term
We consider the Hecke character
$\begin{array}[]{ccc}\mathop{|\delta|}\nolimits:\mathrm{T}({\mathbb{A}})&\rightarrow&{\mathbb{R}}^{\times}_{>0}\\ t&\mapsto&|\delta(t)|\,,\end{array}$
where $\delta$ is the element defined in (14). Note that its restriction to $\mathrm{T}^{\prime}$, seen as a Hecke character of $F$, coincides with the idelic norm on ${\mathbb{I}}_{F}$.
One extends this to a map $|\cdot|:\mathrm{G}({\mathbb{A}})\rightarrow{\mathbb{R}}^{\times}$, by considering a suitable maximal compact subgroup $K_{{\mathbb{A}}}$ of $\mathrm{G}({\mathbb{A}})$ and the associated Iwasawa decomposition
$\mathrm{G}({\mathbb{A}})=K_{{\mathbb{A}}}\mathrm{T}({\mathbb{A}})\rm{U}({\mathbb{A}})$
and letting ||$\cdot$|| ⋅ | be trivial on $\rm{U}({\mathbb{A}})$ and on $K_{{\mathbb{A}}}$.
Now fix an algebraic Hecke character $\phi$ of $\mathrm{T}$, and an element $\Phi\in I^{*}_{\phi}$. Then it is known ([25]) that for $s\in{\mathbb{C}}$ with large enough real part, the infinite sum [^6] Note that we are not using unitary induction. This explains the absence of the half-sum of positive roots in our formulae.
$\sum\limits_{\gamma\in\mathrm{B}({\mathbb{Q}})\backslash\mathrm{G}({\mathbb{Q}})}\Phi(\gamma g)|\gamma g|^{s}$
is absolutely convergent for all $g\in\mathrm{G}({\mathbb{A}})$. Letting $g$ vary, this allows one to define an element
$\mathop{\rm Eis}\nolimits_{\phi,s}(\Phi)\in\mathcal{C}^{\infty}(\mathrm{G}({\mathbb{Q}})\backslash\mathrm{G}({\mathbb{A}}))$
For any $s\in{\mathbb{C}}$ with $\Re(s)>>0$, one thus obtains the Eisenstein operator
(38) $\mathop{\rm Eis}\nolimits_{\phi,s}:I^{*}_{\phi\mathop{|\delta|}\nolimits^{s}}\rightarrow\mathcal{A}(\mathop{{\rm G}}\nolimits({\mathbb{Q}})\backslash\mathop{{\rm G}}\nolimits({\mathbb{A}}))$
As a function of $s$, it admits a meromorphic continuation to the whole of ${\mathbb{C}}$.
Now recall the element $\theta$ of (17). It is well known that the evaluation at $\Psi\in I^{*}_{\phi|\delta|^{s}}$ of the constant term of the operator $\mathop{\rm Eis}\nolimits_{\phi,s}$, formally defined by
$g\mapsto\int\limits_{\rm{U}({\mathbb{Q}})\backslash\rm{U}({\mathbb{A}})}\mathop{\rm Eis}\nolimits_{\phi,s}(\Psi)(ug)\mbox{d}u,\ \ \ g\in\mathrm{G}({\mathbb{A}})$
can be rewritten as
$g\mapsto\Psi(g)+\int\limits_{\rm{U}({\mathbb{A}})}\Psi(\theta\cdot ug)\mbox{d}u,\ \ \ g\in\mathrm{G}({\mathbb{A}})$
and that $\mathop{\rm Eis}\nolimits^{*}_{\phi,s}$ has a pole at $s=0$ if and only if this is the case for the above-described constant term. To study whether such a pole occurs, we introduce the following.
Definition 5.1.
  1. Let $s\in{\mathbb{C}}$ with $\Re(s)>>0$.
  2. (1)
    The intertwining operator
    $T(\phi,s):I^{*}_{\phi\mathop{|\delta|}\nolimits^{s}}\rightarrow I^{*}_{\phi^{-1}\mathop{|\delta|}\nolimits^{-2-s}}$
    is defined, for every $\Psi\in I^{*}_{\phi\mathop{|\delta|}\nolimits^{s}}$ and for every $g\in\mathop{{\rm G}}\nolimits({\mathbb{A}})$, by
    $T(\phi,s)(\Psi(g))=\int\limits_{\rm{U}({\mathbb{A}})}\Psi(\theta\cdot ug)\mbox{d}u\,.$
  3. (2)
    For any place $v$ of ${\mathbb{Q}}$, we define
    $I^{*}_{\phi_{v}\mathop{|\delta_{v}|}\nolimits^{s}}:=\mathop{\rm Ind}\nolimits_{\mathrm{B}({\mathbb{Q}}_{v})}^{\mathrm{G}({\mathbb{Q}}_{v})}\phi_{v}\mathop{|\delta_{v}|}\nolimits^{s}$
    so that $I^{*}_{\phi\mathop{|\delta|}\nolimits^{s}}=\bigotimes\limits_{v}I^{*}_{\phi_{v}\mathop{|\delta_{v}|}\nolimits^{s}}$. Then, the local components $T(\phi_{v},s)$ of the intertwining operator $T(\phi,s)$ are defined, for every $\Psi_{v}\in I^{*}_{\phi_{v}\mathop{|\delta_{v}|}\nolimits^{s}}$ and for every $g_{v}\in\mathop{{\rm G}}\nolimits({\mathbb{Q}}_{v})$, by
    $T(\phi_{v},s)(\Psi_{v}(g_{v}))=\int\limits_{\rm{U}({\mathbb{Q}}_{v})}\Psi_{v}(\theta_{v}\cdot u_{v}g_{v})\mbox{d}u_{v}\,$
    and we have $T(\phi,s)=\bigotimes\limits_{v}T(\phi_{v},s)$.
The pole at $s=0$ of the intertwining operator is controlled by an object of arithmetic nature. To define it, we use the character $\omega_{F|{\mathbb{Q}}}$ introduced in subsection 2.1, and identify Hecke characters of $\mathrm{T}^{\prime}$ to Hecke characters of $F$ according to the conventions of subsection 3.
Definition 5.2.
  1. (1)
    Let $\phi$ be an Hecke character of $\mathrm{T}$, and denote in the same way the Hecke character of $F$ obtained by restricting $\phi$ to $\mathrm{T}^{\prime}$. Denote by $\phi_{{\mathbb{Q}}}$ the restriction of $\phi$ to a Hecke character of ${\mathbb{Q}}$.
    The meromorphic function $c(\phi,s)$ of the complex variable $s\in{\mathbb{C}}$ is defined as
    $c(\phi,s):=\frac{L(\phi,s-1)\cdot L(\phi_{{\mathbb{Q}}}\cdot\omega_{F|{\mathbb{Q}}},2s-2)}{L(\phi,s)\cdot L(\phi_{{\mathbb{Q}}}\cdot\omega_{F|{\mathbb{Q}}},2s-1)}\,.$
    We denote by $\tilde{c}(\phi,s)$ the product of $c(\phi,s)$ with the appropriate $\Gamma$-factors of the occurring $L$-functions.
  2. (2)
    For every place $v$ of ${\mathbb{Q}}$, the local factor $c_{v}(\phi,s)$ of $c(\phi,s)$ is defined as
    $c_{v}(\phi,s):=\frac{L_{p}(\phi,s-1)\cdot L_{p}(\phi_{{\mathbb{Q}}}\cdot\omega_{F|{\mathbb{Q}}},2s-2)}{L_{p}(\phi,s)\cdot L_{p}(\phi_{{\mathbb{Q}}}\cdot\omega_{F|{\mathbb{Q}}},2s-1)}$
    if $v$ corresponds to a finite prime $p$, and as the product of the appropriate $\Gamma$-factors of the occurring $L$-functions, if $v$ is an infinite place. Here $L_{p}(\cdot,\cdot)$ denotes the corresponding factor in the Euler product of the $L$-function under consideration.
Lemma 5.3.
Let $\phi$ be a Hecke character of $\mathrm{T}$. The intertwining operator $T(\phi,s)$ has a pole at $s=0$ if and only if this is the case for $c(\phi,s)$.
Proof.
Let $v=p$ be an odd prime which does not ramify in $F$ and at which $\phi$ is not ramified. Then consider the spherical function $\Psi^{(0)}_{v}\in I^{*}_{\phi_{v}\mathop{|\delta|}\nolimits_{v}^{s}}$ defined for every $g_{v}=b_{v}k_{v}$ with $b_{v}\in\mathrm{B}({\mathbb{Q}}_{v})$ and $k_{v}\in\mathop{{\rm G}}\nolimits({\mathbb{Z}}_{v})$, by $\Psi^{(0)}_{v}(g_{v})=\phi_{v}\mathop{|\delta|}\nolimits_{v}^{s}(b_{v})\,.$ The statement then follows from Lai’s spelling out of Langlands’ constant term formula in our situation ([23, Sec. 3]), saying that
$T(\phi_{v},s)(\Psi^{(0)}_{v})=c_{v}(\phi,s)\cdot\Psi^{(0)}_{v}.$
Hence, we must study the zeroes and poles of the $L$-functions appearing in $c(\phi,s)$. Since the Hecke characters of interest will be those contributing to the description of boundary cohomology of $\mathrm{S}_{K}$, Theorem 4.6 leads us to consider types of the form $w\star\lambda$, where $w$ is an element of the Weyl group of $\mathrm{G}_{F}$ and the characters $w\star\lambda$ have been defined and computed in subsection 2.6. We will only need to consider the case of elements $w$ of length $\geq 2$.
By applying the basic facts about $L$-functions of Hecke characters recalled in subsection 3, we find the following.
Lemma 5.4.
  1. Let $\lambda=(k_{1},k_{2})$ be a character of $\mathrm{T}^{\prime}$ and $w$ an element of the Weyl group of $\mathrm{G}_{F}$ of length $\geq 2$. Fix an Hecke character $\phi$ of $F$ of type $w\star\lambda$.
  2. (1)
    The half-plane of convergence for the Euler product of $L(\phi,s)$ is given by:
    • $\mbox{\emph{Re}}\ s>-\frac{1}{2}-\frac{k_{2}}{2}$ if $w=$ (1 2 3);
    • $\mbox{\emph{Re}}\ s>-\frac{1}{2}-\frac{k_{1}-k_{2}}{2}$ if $w=$ (1 3 2);
    • $\mbox{\emph{Re}}\ s>-1-\frac{k_{1}}{2}$ if $w=$ (1 3).
  3. (2)
    The $L$-function $L(\phi,s)$ is entire, unless $(k_{1},k_{2})=(0,0)$, $w=$ (1 3) and $\phi(\cdot)=|\cdot|_{{\mathbb{I}}_{F}}^{2}$. In this case, it has a unique first-order pole at $s=-1$.
Moreover, taking $\phi$ as above, its restriction $\phi_{{\mathbb{Q}}}$ to ${\mathbb{I}}_{{\mathbb{Q}}}$ has weight $2w(\phi)$. The same then holds for $\phi_{{\mathbb{Q}}}\cdot\omega_{F|{\mathbb{Q}}}$. Hence, we obtain the following.
Lemma 5.5.
  1. Let $\lambda=(k_{1},k_{2})$ be a character of $\mathrm{T}^{\prime}$ and $w$ an element of the Weyl group of $\mathrm{G}_{F}$ of length $\geq 2$. Fix an Hecke character $\phi$ of $F$ of type $w\star\lambda$.
  2. (1)
    The half-plane of convergence for the Euler product of $L(\phi_{{\mathbb{Q}}}\cdot\omega_{F|{\mathbb{Q}}},s)$ is given by:
    • $\mbox{\emph{Re}}\ s>-2-k_{2}$ if $w=$ (1 2 3);
    • $\mbox{\emph{Re}}\ s>-2-(k_{1}-k_{2})$ if $w=$ (1 3 2);
    • $\mbox{\emph{Re}}\ s>-3-k_{1}$ if $w=$ (1 3).
  3. (2)
    The $L$-function $L(\phi\cdot\omega_{F|{\mathbb{Q}}},s)$ is entire, unless
    $k_{2}=0,\ w=\mbox{\emph{(1\ 2\ 3)}}\ \mbox{and}\ \phi_{{\mathbb{Q}}}\cdot\omega_{F|{\mathbb{Q}}}(\cdot)=|\cdot|_{{\mathbb{I}}_{\mathbb{Q}}}^{3}\,,$
    or
    $k_{1}=k_{2},\ w=\mbox{\emph{(1\ 3\ 2)}}\ \mbox{and}\ \phi_{{\mathbb{Q}}}\cdot\omega_{F|{\mathbb{Q}}}(\cdot)=|\cdot|_{{\mathbb{I}}_{\mathbb{Q}}}^{3}\,.$
    In each of the above two cases, it has a unique first-order pole at $s=-2$.
From the previous two lemmas, we get immediately the following corollary.
Corollary 5.6.
[…]
If $L(\phi,-1)=0$, then $c(\phi,s)$ has no pole at $s=0$.
5.3. Automorphic description of Eisenstein cohomology
In order to describe the main result of this section, we need to consider a modification of the intertwining operators $T(\phi,s)$ introduced in the previous subsection.
Definition 5.7.
Let $\lambda$ be a character of $\mathrm{T}^{\prime}$ and $w$ an element of $W$ of length $\geq 2$. Let $\phi$ be an Hecke character of $F$ of type $w\star\lambda$. Let $s\in{\mathbb{C}}$ belong to the half-plane of convergence of the Euler product of $L(\phi,s)$. For every place $v$, we define
$T^{\mathop{\rm loc}\nolimits}(\phi_{v},s):=c_{v}(\phi,s)^{-1}\cdot T(\phi_{v},s)\,.$
We then put
$T^{\mathop{\rm loc}\nolimits}(\phi,s):=\bigotimes\limits_{v}T^{\mathop{\rm loc}\nolimits}(\phi_{v},s)\,.$
Remark 5.8.
  1. (1)
    If $s$ satisfies the hypotheses, i.e. $\mbox{Re}\ s$ satisfies the bounds of Lemma 5.4.(1), then both the operator $T(\phi_{v},s)$ and the quantity $c_{v}(\phi,s)$ are defined and non-zero (cfr. [17, Lemma 2.3.1]).
  2. (2)
    The operator $T^{\mathop{\rm loc}\nolimits}(\phi,s)$ is also defined, and non-zero, for $s=0$.
For the rest of this section, we suppose that $\lambda$, $w$ and $\phi$ are as in the definition above.
Definition 5.9.
The operator $\mathop{\rm Eis}\nolimits_{\phi}:I_{\phi}^{*}\rightarrow\mathcal{A}(\mathop{{\rm G}}\nolimits({\mathbb{Q}})\backslash\mathop{{\rm G}}\nolimits({\mathbb{A}}))$ is defined as
  1. (1)
    the evaluation $\mathop{\rm Eis}\nolimits_{\phi,0}$ at $s=0$ of the Eisenstein operator $\mathop{\rm Eis}\nolimits_{\phi,s}$ of (38), if $c(\phi,s)$ does not have a pole at $s=0$;
  2. (2)
    the residue at $s=0$ of the operator $s\cdot\mathop{\rm Eis}\nolimits_{\phi,s}$, if $c(\phi,s)$ has a pole at $s=0$.
Definition 5.10.
Denote by $\Phi_{\mbox{\tiny{np}}}$ the set of those $\phi$’s such that $c(\phi,s)$ does not have a pole at $s=0$, and by $\Phi_{\mbox{\tiny{p}}}$ the set of those $\phi$’s such that $c(\phi,s)$ has a pole at $s=0$.
  1. (1)
    For each $\phi\in\Phi_{\mbox{\tiny{np}}}$, we denote by $\mathop{\rm Eis}\nolimits_{\phi}$ the operator induced in $(\mathfrak{g},\tilde{K}_{\infty})$-cohomology by the the operator $\mathop{\rm Eis}\nolimits_{\phi}$ of Definition 5.9.(1).
  2. (2)
    For each $\phi\in\Phi_{\mbox{\tiny{p}}}$, we denote by $\mathop{\rm Eis}\nolimits^{\prime}_{\phi}$ the restriction to $I^{\prime}_{\phi}$ of the operator $\mathop{\rm Eis}\nolimits_{\phi}$ induced in $(\mathfrak{g},\tilde{K}_{\infty})$-cohomology by the operator $\mathop{\rm Eis}\nolimits_{\phi}$ of Definition 5.9.(2).
Definition 5.11.
  1. (1)
    We denote $I^{\prime}_{\phi}$ the submodule of $I_{\phi}$ generated by those tensors, whose component at $v$ belongs to the kernel of $T^{\mathop{\rm loc}\nolimits}(\phi_{v},0)$ for at least one finite place $v$.
  2. (2)
    For any finite place $v$, we denote $\tilde{J}_{\phi_{v}}:=\mbox{Im}\ T^{\mathop{\rm loc}\nolimits}(\phi_{v},0)$, and
    (39) $J_{\phi}:=\bigotimes_{v}\tilde{J}_{\phi,v}$
Remark 5.12.
As explained in [17, p. 583], the results of [20] imply that if $\phi\in\Phi_{p}$, then the spaces $I^{\prime}_{\phi}$, resp. $J_{\phi}$ are non-trivial, proper subspaces of the domain, resp. of the codomain of $T^{\mathop{\rm loc}\nolimits}(\phi,0)$.
We are now ready to state the main result of Harder’s paper [17].
Theorem 5.13.
[17, Thm. 2] Let $r$ be the restriction morphism of (36). Then, $r_{{\mathbb{C}}}\circ(\oplus_{\phi\in\Phi_{\mbox{\tiny{np}}}}\mathop{\rm Eis}\nolimits_{\phi}\oplus_{\phi\in\Phi_{\mbox{\tiny{p}}}}\mathop{\rm Eis}\nolimits^{\prime}_{\phi})$ induces an isomorphism
(40) $\mathop{\rm Im}\nolimits(r)\simeq\bigoplus\limits_{w\in{\mathcal{W}}}\bigoplus_{\begin{subarray}{c}\mbox{\emph{\tiny{type}}}(\phi)=w\star\lambda\\ \ell(w)\geq 2\\ \phi\in\Phi_{\mbox{\tiny{np}}}\end{subarray}}I_{\phi}\bigoplus_{\begin{subarray}{c}\mbox{\emph{\tiny{type}}}(\phi)=w\star\lambda\\ \ell(w)\geq 2\\ \phi\in\Phi_{\mbox{\tiny{p}}}\end{subarray}}I^{\prime}_{\phi}\oplus J_{\phi}$
Remark 5.14.
  1. (1)
    The case $\phi\in\Phi_{\mbox{\tiny{np}}}$ of the above theorem is proven in [17, pp. 581-583], and the case $\phi\in\Phi_{\mbox{\tiny{p}}}$ in [17, pp. 583-584].
  2. (2)
    In the above theorem, the spaces $I_{\phi}$ and $I^{\prime}_{\phi}$ contribute to cohomology in degrees corresponding to the length of the corresponding Weyl element, hence degree 2 (in which case the space $J_{\phi}$ contributes to cohomology in degree 1) or 3 (in which case the space $J_{\phi}$ contributes to cohomology in degree 0).
  3. (3)
    The map in the statement of the above theorem induces an isomorphism between the extensions of scalars to ${\mathbb{C}}$ of the spaces under consideration. One then obtains the desired isomorphism by passing to suitable $\overline{{\mathbb{Q}}}$-structures.
For $i\in\{0,\dots,3\}$, we will denote
(41) $H^{i}_{\mathop{\rm Eis}\nolimits}(\mathrm{S}({\mathbb{C}}),V_{\lambda}):=\mbox{Im}(H^{i}(\mathrm{S}({\mathbb{C}}),V_{\lambda})\rightarrow\partial H^{i}(\mathrm{S}({\mathbb{C}}),V_{\lambda}))$
where the image is taken with respect to the restriction morphism. We call this space Eisenstein cohomology, and its counterpart at finite level is obtained by taking $K$-invariants,
$H^{\bullet}_{\mathop{\rm Eis}\nolimits}(\mathrm{S}_{K}({\mathbb{C}}),V_{\lambda})=H^{\bullet}_{\mathop{\rm Eis}\nolimits}(\mathrm{S}({\mathbb{C}}),V_{\lambda})^{K}$
With this notation, the first concrete consequence of Theorem 5.13 that we will use is the following.
Corollary 5.15.
Let $k$ be a positive integer and let $\phi$ be a Hecke character of $F$ of infinity type $(k,-k-3)$. Suppose $\phi_{{\mathbb{Q}}}\cdot\omega_{F|{\mathbb{Q}}}(\cdot)=|\cdot|_{{\mathbb{I}}_{\mathbb{Q}}}^{3}$.
If $L(\phi,-1)=0$, then there exists a canonical monomorphism of $\mathrm{G}({\mathbb{A}}_{f})$-modules
$I_{\phi}\hookrightarrow H^{2}_{\mathop{\rm Eis}\nolimits}(\mathrm{S}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})$
For any compact open $K\subset\mathrm{G}({\mathbb{A}}_{f})$ which is small enough, it endows $I^{K}_{\phi}$ with a pure Hodge structure, direct sum of 1-dimensional pure $\overline{{\mathbb{Q}}}$-Hodge structures
$I_{k,-k-3}\otimes_{F}\overline{{\mathbb{Q}}}$
of type $(1,k+2)$.
Proof.
[…]
In order to state the second consequence of Theorem 5.13 which will be of interest for us, denote, for $i\in\{0,\dots,3\}$
(42) $H^{i}_{c,\mathop{\rm Eis}\nolimits}(\mathrm{S}({\mathbb{C}}),V_{\lambda}):=\mbox{Coker}(H^{i}(\mathrm{S}({\mathbb{C}}),V_{\lambda})\rightarrow\partial H^{i}(\mathrm{S}({\mathbb{C}}),V_{\lambda}))$
where again the cokernel is taken with respect to the restriction morphism. Moreover, for $\phi$ a Hecke character of $F$ of type $\lambda$, define $\theta\phi$ by
$\theta\phi:=\theta(\phi|\delta|)|\delta|^{-1}$
Corollary 5.16.
Let $k$ and $\phi$ be as in the Corollary above. Then there exists a canonical monomorphism of $\mathrm{G}({\mathbb{A}}_{f})$-modules
$I_{\theta\phi}\hookrightarrow H^{1}_{c,\mathop{\rm Eis}\nolimits}(\mathrm{S}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})$
exhibiting $I_{\theta\phi}$ as a direct summand of the space on the right.
For any compact open $K\subset\mathrm{G}({\mathbb{A}}_{f})$ which is small enough, it endows $I^{K}_{\theta\phi}$ with a pure Hodge structure, direct summand of the one on $H^{1}_{c,\mathop{\rm Eis}\nolimits}(\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})$. It is the direct sum of 1-dimensional pure $\overline{{\mathbb{Q}}}$-Hodge structures
$I_{k+1,-k-2}\otimes_{F}\overline{{\mathbb{Q}}}$
of type $(0,k+1)$
Proof.
[…]
Notation 5.17.
Fix $\phi$ as in the above Corollary, and $K$ small enough. We will denote the pure Hodge structure having $I^{K}_{\phi}$ as underlying vector space by
(43) $\partial H^{2}(\phi)$
and the one having $I^{K}_{\theta\phi}$ as underlying vector space by
(44) $\partial H^{1}(\theta\phi)$
Remark 5.18.
Suppose $\phi_{{\mathbb{Q}}}$ satisfies the same assumption as in Corollary 5.15, but $\phi$ is of infinity type $(-k-3,k)$. Then, using the local system $V^{\vee}_{k}$ instead of $V_{k}$, and employing Corollary 4.4, the same proof shows the existence of a sub-Hodge structure $\partial H^{2}(\phi)$ of degree-2 boundary cohomology, of Hodge type $(k+1,0)$ and of underlying vector space $I^{K}_{\phi}$.
6. A subspace of interior cohomology
The aim of this section is to construct specific sub-Hodge structures $H^{2}(\phi)_{!}$ of interior cohomology of Picard surfaces. These subspaces are associated with Hecke characters $\phi$ of the imaginary quadratic field $F$ satisfying the assumptions of our main theorem, Theorem 1, and provide indeed the last ingredient which will be needed to prove it in the next section.
6.1. The setup
[…]
We will consider those $\phi$’s such that the restriction $\phi_{{\mathbb{Q}}}$ of $\phi$ to ${\mathbb{Q}}$ satisfies
(45) $\phi_{{\mathbb{Q}}}=\omega_{F|{\mathbb{Q}}}|\cdot|^{3}_{{\mathbb{I}}_{{\mathbb{Q}}}}$
Remark 6.1.
Let us discuss the existence of Hecke characters verifying the above assumption. Observe that an Hecke character $\phi$ of $F$ infinity type $(k,-k-3)$ can be written as
(46) $\phi=|\delta|^{\frac{3}{2}}\cdot\phi_{u}$
where $\phi_{u}$ is a unitary Hecke character. Then we have an equality of $\mathrm{G}({\mathbb{A}}_{f})$-representations
(47) $\mathop{\rm Ind}\nolimits_{\mathrm{B}({\mathbb{A}}_{f})}^{\mathrm{G}({\mathbb{A}}_{f})}\phi=\mathop{\rm IndUn}\nolimits_{\mathrm{B}({\mathbb{A}}_{f})}^{\mathrm{G}({\mathbb{A}}_{f})}\phi_{u}\cdot|\delta|^{\frac{1}{2}}$
where $\mathop{\rm IndUn}\nolimits$ denotes the unitary induction functor. Moreover, we have $\phi_{u,\infty}(z)=z^{-k-\frac{3}{2}}\overline{z}^{k+\frac{3}{2}}$ for $z\in{\mathbb{C}}^{\times}$.
Now, the proof of [4, Lemma 6.9.2] shows: $(a)$ that condition (45) is equivalent to asking that $\phi_{u}=\overline{\phi_{u}}^{-1}$; $(b)$ that such $\phi_{u}$’s verifying the latter condition do exist.
The second condition that we want to enforce on our $\phi$ is that
(48) $L(\phi,-1)=0$
Under (45), the intertwining operator $T(\phi,s)$ may have a pole at $s=0$, but when (48) holds, this is not the case; hence, $T(\phi,0)$ is well defined.
Definition 6.2.
Let $k$ be a positive integer and $\phi$ be an Hecke character of $F$ of infinity type $(k,-k-3)$. If $\phi$ satisfies (45) and (48), then we denote the Langlands quotient of $I^{*}_{\phi}$ by
$J^{*}_{\phi}:=\mbox{Im}(T(\phi,0))$
We denote $J_{\phi,v}$ its component at a finite place $v$, and $J_{\phi,f}$ the restricted product of the $J_{\phi,v}$’s over $v$.
Remark 6.3.
  1. (1)
    By the results of [20] (see also [29, 12.2]), asking that $\phi$ verifies (45) implies that $J^{*}_{\phi}$ is non-trivial.
  2. (2)
    Recall the space $J_{\phi}$ introduced in (39). By construction, we have $J_{\phi,{\mathbb{C}}}\simeq J_{\phi,f}$.
Observe that if we suppose that the sign $\epsilon:=\epsilon(-1)$ of the functional equation of $\phi$ satisfies
(49) $\epsilon=-1$
then (48) holds. Now, consider the space $\mathrm{S}$ introduced in subsection 4.4. Recall that for any local system $V_{\lambda}$, the cuspidal cohomology
$H^{\bullet}_{\mathop{\rm\tiny{cusp}}\nolimits}(S({\mathbb{C}}),V_{\lambda,{\mathbb{C}}})$
is the sub-$\mathrm{G}({\mathbb{A}}_{f})$-module of interior cohomology $H_{!}^{\bullet}(S({\mathbb{C}}),V_{\lambda,{\mathbb{C}}})$ (cfr. (32)), defined as the direct sum of all the irreducible sub-$\mathrm{G}({\mathbb{A}}_{f})$-modules of $H^{\bullet}_{!}(S({\mathbb{C}}),V_{\lambda})$ isomorphic to the non-archimedean component of some cuspidal automorphic representation of $\mathrm{G}({\mathbb{A}})$. If we denote the set of these isomorphism classes by $\mathcal{C}_{\mathop{\rm\tiny{cusp}}\nolimits}$, then by multiplicity one for the discrete spectrum of $\mathrm{G}({\mathbb{A}})$ ([29, Thm. 13.3.1]), cuspidal cohomology decomposes as
(50) $H^{\bullet}_{\mathop{\rm\tiny{cusp}}\nolimits}(S({\mathbb{C}}),V_{\lambda,{\mathbb{C}}})\simeq\bigoplus\limits_{\begin{subarray}{c}\pi=\pi_{f}\otimes\pi_{\infty}\\ \mbox{\tiny{s. t.}}\ \pi_{f}\in\mathcal{C}_{\mathop{\rm\tiny{cusp}}\nolimits}\end{subarray}}\pi_{f}\otimes H^{\bullet}({\mathfrak{g}},\mathop{\tilde{\rm K}_{\infty}}\nolimits,\pi_{\infty}\otimes V_{\lambda,{\mathbb{C}}})$
In [18, p. 99], the following has been observed. It is a consequence of [29, Thms. 13.3.6, 13.3.7], as corrected in [31, Thm. 1.1], that under (49), there exists a $\mathrm{G}({\mathbb{A}}_{f})$-submodule $H^{2}(\pi(\phi)_{f})$ of $H^{2}_{\mathop{\rm\tiny{cusp}}\nolimits}(S({\mathbb{C}}),V_{k})$, whose local components are isomorphic to $J^{*}_{\phi,v}$ for infinitely many places $v$.
6.2. An application of Rogawski’s results
Given the Hecke character $\phi$, we are now going to construct the $\mathrm{G}({\mathbb{A}}_{f})$-submodule $H^{2}(\pi(\phi)_{f})$ mentioned above, by constructing a representation of $\mathrm{G}({\mathbb{A}})$ which we will show to be cuspidal and cohomological as an application of Rogawski’s results.
For this, we fix a choice of a Hecke character $\mu$ of $F$ whose restriction to ${\mathbb{Q}}$ is $\omega_{F|{\mathbb{Q}}}(\cdot)$. Moreover, we choose a couple $(\rho,\rho^{\prime})$ of characters on the norm-one subgroup $C_{F}$ of the idèle class group of $F$ as follows: define $\rho_{1}$ to be the restriction of $\mu\cdot\phi_{u}$ to $C_{F}$, and $\rho^{\prime}$ to be trivial. To such a couple, one attaches a character $\rho$ on $\mathrm{U}_{2}\times\mathrm{U}_{1}$ as in [31, p. 395].
Definition 6.4.
Let $\phi$ be a Hecke character of $F$ satisfying
$\phi_{{\mathbb{Q}}}=\omega_{F|{\mathbb{Q}}}|\cdot|^{3}_{{\mathbb{I}}_{{\mathbb{Q}}}}$
For $v$ a finite place of ${\mathbb{Q}}$, define the $\mathrm{G}({\mathbb{Q}}_{v})$-representation $\pi^{n}(\rho_{v})$ in the following way:
  1. (1)
    is $v$ is inert or ramified in $F$, we define it as the Langlands quotient $J_{\phi,v}$ of $I_{\phi,v}$, (non-trivial under our hypothesis on $\phi$, by Rmk. 6.3);
  2. (2)
    if $v$ splits into places $w_{1}$ and $w_{2}$ in $F$, identifying $\mathrm{G}({\mathbb{Q}}_{v})\simeq\mathop{\rm GL}\nolimits_{3}(F_{w_{1}})\times F^{\times}_{w_{1}}$, and seeing $\phi_{v}$ as a character of $(F^{\times}_{w_{1}})^{4}$, we define it as the representation induced from the standard $(2,1)$ parabolic of $\mathop{\rm GL}\nolimits_{3}(F_{w_{1}})$ by the character
    $\left(\left(\begin{array}[]{cc}M&*\\ &\lambda\end{array}\right),f\right)\mapsto\phi_{v}(\det(M),1,\lambda,f)$
Remark 6.5.
For any finite place $v$, the unitarily induced representation called $\mathop{\rm ind}\nolimits(\eta_{v})$ in [31, p. 395] is nothing but our $I_{\phi_{v}}$, or more precisely, its restriction to a representation of $\mathrm{G}^{\nu=1}$. This follows by observing that in the notation of (47), the character called $\phi$ in loc. cit. becomes $\phi_{u}$. This shows that for $v$ inert or ramified, our notation $\pi^{n}(\rho_{v})$ is coherent with Rogawski’s one in [31, p. 395]. The same holds, by definition, in the case of split $v$ ([31, p. 396]).
Definition 6.6.
The discrete series representation $\pi(\phi)_{\infty}$ of $\mathrm{G}({\mathbb{R}})$ is defined as the representation $\pi^{s}(\rho_{\infty})$ denoted by $\pi^{s}(\rho)$ in [30, p. 77, first line].
Remark 6.7.
Let $\lambda=(k_{1},k_{2},c,r)$ be a character of $\mathrm{T}_{F}$. By [30, 3.2], the representation $\pi^{s}(\rho)$ is $V_{\lambda}$-cohomological in degree 2 when either $k_{2}=0$ or $k_{1}=k_{2}$. In the first case, it coincides with the discrete series representation denoted by $\pi^{+}$ in loc. cit.; in the second case, it coincides with the one denoted by $\pi^{-}$.
To see how this follows from loc. cit., observe that in our notation, the restriction to $\mathrm{G}^{\nu=1}$ of the representation $V_{\lambda}$ is identified with a triple $(k_{1},k_{2},c)$, whereas in [30, p. 79] it is identified with a triple $(m,r,n)$. Now, under the change of variables between the two parametrizations, one sees that $k_{2}=0$ if and only if $r-n=1$, and that $k_{1}=k_{2}$ if and only if $m-r=1$.
Proposition 6.8.
Let $k$ be a positive integer and $\phi$ an Hecke character of $F$ of type $(k,-k-3)$, satisfying
$\phi_{{\mathbb{Q}}}=\omega_{F|{\mathbb{Q}}}|\cdot|^{3}_{{\mathbb{I}}_{{\mathbb{Q}}}}$
and
$\epsilon(\phi)=-1$
Let the $\mathrm{G}({\mathbb{A}}_{f})$-representation $\pi_{f}(\phi)$ and the $\mathrm{G}({\mathbb{A}})$-representation $\pi(\phi)$ be defined by
$\pi(\phi)_{f}$ $:=\bigotimes\limits_{v}\pi^{n}(\rho_{v})\,,$
(51) $\pi(\phi)$ $:=\pi(\phi)_{f}\otimes\pi(\phi)_{\infty}$
Then
  1. (1)
    The representation $\pi(\phi)$ is automorphic and has multiplicity 1 in the cuspidal spectrum of $\mathrm{G}$.
  2. (2)
    There exist a $\overline{{\mathbb{Q}}}$-model of $\pi_{f}(\phi)$, denoted by the same symbol, a 1-dimensional $\overline{{\mathbb{Q}}}$-Hodge structure $H^{2}(\phi)_{!}$ of type $(k+2,0)$, and, for any neat open compact $K$ small enough, a submodule under the Hecke algebra of level $K$
    $H^{2}(\pi(\phi)_{f})\hookrightarrow H^{2}_{!}(\mathrm{S}_{K}({\mathbb{C}}),V_{k}\otimes_{F}\bar{{\mathbb{Q}}})$
    such that
    (52) $H^{2}(\pi(\phi)_{f})\simeq\pi(\phi)_{f}^{K}\otimes H^{2}(\phi)_{!}$
    as a Hecke module and as a Hodge structure.
Proof.
  1. (1)
    The automorphic representation $\pi(\phi)$ belongs to the $A$-packet $\Pi(\rho)$ of [31, p. 396], and by construction, the number $n(\pi)$ of loc. cit. of places $v$ where $\pi(\phi)_{v}$ is isomorphic to $\pi^{s}(\rho_{v})$ is equal to 1 (since this happens only for $v=\infty$). Now write $\phi_{R}$ for the character denoted $\phi$ in loc. cit.. By Remark 6.5, we have $\phi_{R}=\phi^{(1)}_{u}$, so that by (46)
    $L\left(\phi_{R},\frac{1}{2}\right)=L\left(\phi\cdot|\cdot|_{L}^{-\frac{3}{2}},\frac{1}{2}\right)=L\left(\phi,-1\right)$
    which shows that our hypothesis $\epsilon(\phi)=-1$ is precisely equivalent to $\epsilon(\frac{1}{2},\phi_{R})=-1$. By [31, Thm. 1.1], we obtain that the multiplicity of $\pi(\phi)$ in the discrete spectrum of $\mathrm{G}$ is 1. By [29, Thm. 13.3.6 (a)], it is actually a representation appearing in the cuspidal spectrum.
  2. (2)
    Consider the decomposition (50) of $V_{k}$-valued cuspidal cohomology of $\mathrm{S}$. Upon choosing a $\overline{{\mathbb{Q}}}$-model for each submodule $\pi_{f}$, denoted by the same symbol, we get a $\overline{{\mathbb{Q}}}$-model for cuspidal cohomology, described as
    (53) $H^{\bullet}_{\mathop{\rm\tiny{cusp}}\nolimits}(\mathrm{S}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}}):=\bigoplus\limits_{\begin{subarray}{c}\pi=\pi_{f}\otimes\pi_{\infty}\\ \mbox{\tiny{s. t.}}\ \pi_{f}\in\mathcal{C}_{\mathop{\rm\tiny{cusp}}\nolimits}\end{subarray}}\pi_{f}\otimes\mathop{\rm Hom}\nolimits_{\mathrm{G}({\mathbb{A}}_{f})}(\pi_{f},H^{\bullet}_{!}(\mathrm{S}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}}))$
    where each space $\mathop{\rm Hom}\nolimits_{\mathrm{G}({\mathbb{A}}_{f})}(\pi_{f},H^{\bullet}_{!}(\mathrm{S}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}}))$ is endowed with a canonical $\overline{{\mathbb{Q}}}$-Hodge structure. Now consider the submodule corresponding to $\pi=\pi(\phi)$. By Remark 6.7, $\pi(\phi)_{\infty}$ is $V_{k}$-cohomological in degree 2. By the comparison isomorphism between singular interior cohomology and étale interior cohomology, the dimension of the Hodge structure
    (54) $H^{2}(\phi)_{!}:=\mathop{\rm Hom}\nolimits_{\mathrm{G}({\mathbb{A}}_{f})}(\pi(\phi)_{f},H^{2}_{!}(\mathrm{S}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}}))$
    is the same as the dimension of the space $V^{2}(\pi_{f}))$ of Case 5 of [30, pp. 91-92], which according to loc. cit. is equal to 1, by our choice of the archimedean component of $\pi(\phi)$. We obtain then the desired Hecke submodule. Finally, to get the assertion on the type of $H^{2}(\phi)_{!}$, we apply the following two facts:
    (a) if $r$ is the weight of the pure Hodge structure on $V_{\lambda}$ (so that, by (22), $r=k$ in our case), then $H^{2}_{!}(\mathrm{S}({\mathbb{C}}),V_{\lambda})$ has weight $r+2$. Its Hodge substructure $H^{2}(\pi(\phi)_{f})$ has thus the same weight;
    (b) the Hodge types of $H^{2}(\pi(\phi)_{f})$ are those $(p,q)$’s appearing in the $(p,q)$-decomposition of the space $H^{2}({\mathfrak{g}},\mathop{\tilde{\rm K}_{\infty}}\nolimits,\pi(\phi)_{\infty}\otimes V_{k,{\mathbb{C}}})$, which is in turn determined by the $(p,q)$-decomposition of the discrete series representation $\pi(\phi)_{\infty}$. Now, by Remark 6.7 the discrete series representation $\pi(\phi)_{\infty}$ coincides with $\pi^{+}$, which is holomorphic by ([7, 1.3]).
Remark 6.9.
Suppose the hypotheses of Proposition 6.8 hold. For any small enough level $K$, the simple, level $K$ Hecke module $\pi^{K}(\phi)_{f}$ corresponds to a minimal ideal in the level $K$ Hecke algebra. Take an idempotent generator $e_{\pi(\phi)}$ of this ideal. By a standard argument (cfr. for example [35, Prop. 5.4]) we see that the Hodge structure $H^{2}(\phi)_{!}$ is isomorphic to the image of $H_{!}^{2}(\mathrm{S}_{K}({\mathbb{C}}),V_{k})$ under $e_{\pi(\phi)}$. Thus, we will be able to consider $H^{2}(\phi)_{!}$ both as a subobject and as a quotient of the Hodge structure on $H_{!}^{2}(\mathrm{S}_{K}({\mathbb{C}}),V_{k})$.
Remark 6.10.
Suppose $\phi_{{\mathbb{Q}}}$ satisfies the assumption of Prop. 6.8, but $\phi$ is of infinity type $(-k-3,k)$. The same proof, but employing the local system $V^{\vee}_{k}$ instead of $V_{k}$, shows the existence of a Hodge structure $H^{2}(\phi)_{!}$ of type $(0,k+2)$ satisfying all the other requirements in the statement. This follows from the computation of the highest weight of $V^{\vee}_{k}$ of Rmk. 2.2 and from the fact that using the latter local system, by Rmk. 6.7, the infinite component $\pi(\phi)_{\infty}$ is isomorphic to the discrete representation $\pi^{-}$, which is anti-holomorphic by ([7, 1.3]).
7. Extensions and biextensions associated to Hecke characters
In this last section, we keep on considering Hecke characters $\phi$ of $F$ satisfying the hypotheses of the previous section, which will be recalled below. In particular, their $L$-function is supposed to vanish at the central point, and as a consequence, as discussed in 3.3, the Bloch-Beilinson conjectures predict the existence of a non-trivial extension of Hodge structures $E_{\phi}$ of geometric origin.
Our first aim is to put together all the ingredients introduced so far, in order to construct candidates $E_{\phi}(\Psi)$ for $E_{\phi}$, by means of the cohomology of a local system on a Picard modular surface $\mathrm{S}_{K}$. They depend on the choice of a cohomology class $\Psi$ in a suitable subspace of Eisenstein cohomology of $\mathrm{S}_{K}$. This will prove Theorem 1 in the introduction.
In the second and last part of this section, we will show how such $E_{\phi}(\Psi)$’s sit naturally in certain biextensions of Hodge structures. The relevance of this construction for our program of proving non-triviality of (some) $E_{\phi}(\Psi)$ has been discussed in the introduction.
7.1. Construction of extensions
For a positive integer $k$ and a neat level $K\subset\mathrm{G}({\mathbb{A}}_{f})$, we look at the local system $V_{k}$ over the Picard modular surface $\mathrm{S}_{K}$. Consider its interior cohomology, defined in (32), and its Eisenstein cohomology, defined in (41). The extension of scalars to $\overline{{\mathbb{Q}}}$ of the long exact sequence (28) gives rise to a short exact sequence
(55) $0\rightarrow H_{!}^{2}(\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})\rightarrow H^{2}(\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})\rightarrow H_{\mathop{\rm Eis}\nolimits}^{2}(\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})\rightarrow 0$
In order to state our main theorem, recall that to a Hecke character $\phi$ satisfying the hypotheses formulated below, we have attached pure Hodge structures $\partial H^{2}(\phi)$ and $H(\phi)_{!}$, respectively in (43) and (54).
Theorem 7.1.
Let $k$ be a positive integer and $\phi$ an Hecke character of $F$ of type $(k,-k-3)$, satisfying
$\phi_{{\mathbb{Q}}}=\omega_{F|{\mathbb{Q}}}|\cdot|^{3}_{{\mathbb{I}}_{{\mathbb{Q}}}}$
and
$\epsilon(\phi)=-1$
  1. (1)
    For any level $K$ small enough, the exact sequence (55) has a subquotient short exact sequence of Hodge structures
    (56) $0\rightarrow H^{2}(\phi)_{!}\rightarrow\tilde{E}_{0}\rightarrow\partial H^{2}(\phi)\rightarrow 0$
  2. (2)
    Let $m(\phi,K)$ be the dimension of $I^{K}_{\phi}$. Then the short exact sequence (56) induces an extension of $\overline{{\mathbb{Q}}}$-Hodge structures
    (57) $0\rightarrow H_{\phi}(-1)\rightarrow E_{0}\rightarrow\mathds{1}^{\oplus m(\phi,K)}\rightarrow 0$
    where $H_{\phi}$ is the 1-dimensional $\overline{{\mathbb{Q}}}$-Hodge structure of type $(k,-(k+3))$ attached to $\phi$.
Proof.
  1. (1)
    Since $\epsilon(\phi)=-1$, we have $L(\phi,-1)=0$. Thus, by Corollary 5.15, the pure Hodge structure $\partial H^{2}(\phi)$ is a sub-Hodge structure of $H^{2}_{\mathop{\rm Eis}\nolimits}(\mathrm{S}_{K}({\mathbb{C}}),V_{k})$. By Proposition 6.8 and Remark 6.9, the pure Hodge structure $H^{2}(\phi)_{!}$ is a quotient Hodge structure of $H^{2}_{!}(\mathrm{S}_{K}({\mathbb{C}}),V_{k})$. Hence, the desired extension is obtained from the short exact sequence (55) by taking pullback via the inclusion $\partial H^{2}(\phi)\hookrightarrow H^{2}_{\mathop{\rm Eis}\nolimits}(\mathrm{S}_{K}({\mathbb{C}}),V_{k})$ and pushout via the projection $H^{2}_{!}(\mathrm{S}_{K}({\mathbb{C}}),V_{k})\twoheadrightarrow H^{2}(\phi)_{!}$.
  2. (2)
    By Corollary 5.15, the pure Hodge structure $\partial H^{2}(\phi)$ is a direct sum of copies of the 1-dimensional $\overline{{\mathbb{Q}}}$-Hodge structure $I_{k,-k-3,\overline{{\mathbb{Q}}}}$. By tensoring the extension of the previous point by the dual of $I_{k,-k-3,\overline{{\mathbb{Q}}}}$, we get an extension
    $0\rightarrow H^{2}(\phi)_{!}\otimes I^{\vee}_{k,-k-3,\overline{{\mathbb{Q}}}}\rightarrow E_{0}\rightarrow\mathds{1}^{\oplus m(\phi,K)}\rightarrow 0$
    Since the type of $H^{2}(\phi)_{!}$ is $(k+2,0)$ and the type of $I^{\vee}_{k,-k-3,\overline{{\mathbb{Q}}}}$ is $(1,k+2)$, the type of the 1-dimensional $\overline{{\mathbb{Q}}}$-Hodge structure $H^{2}(\phi)_{!}\otimes I^{\vee}_{k,-k-3,\overline{{\mathbb{Q}}}}$ is $(k+1,-k-2)$. This tells us that as a $\overline{{\mathbb{Q}}}$-Hodge structure, $H^{2}(\phi)_{!}\otimes I^{\vee}_{k,-k-3,\overline{{\mathbb{Q}}}}$ is indeed isomorphic to $H_{\phi}(-1)$.
Corollary 7.2.
Let $k$ and $\phi$ be as in Theorem 57. The choice of an element $\Psi$ of $I_{\phi}^{K}$ provides, by pullback of the extension (57), a 1-extension $E_{\phi}(\Psi)$ of $\mathds{1}$ by $H_{\phi}(-1)$ in the category $\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}}$. This defines a morphism
(58) $I^{K}_{\phi}\rightarrow\mathop{\rm Ext}\nolimits^{1}_{\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}}}(\mathds{1},H_{\phi}(-1))$
Using Remarks 5.18 and 6.10, one can run the same proof as the one of Theorem 57, but employing $V^{\vee}_{k}$ instead of $V_{k}$, and deduce the following variant of the above Corollary.
Corollary 7.3.
Let $k$ and $\phi_{{\mathbb{Q}}}$ be as in Theorem 57, but suppose that $\phi$ is of infinity type $(-k-3,k)$. The choice of an element $\Psi$ of $I_{\phi}^{K}$ provides a 1-extension $E_{\phi}(\Psi)$ of $\mathds{1}$ by $H_{\phi}(-1)$ in the category $\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}}$. This defines a morphism
(59) $I^{K}_{\phi}\rightarrow\mathop{\rm Ext}\nolimits^{1}_{\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}}}(\mathds{1},H_{\phi}(-1))$
For the needs of the next subsection, we also need to construct a dual extension, in the following way. The extension of scalars to $\overline{{\mathbb{Q}}}$ of the long exact sequence (28) gives rise to a short exact sequence
(60) $\partial H^{1}(\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})\rightarrow H_{c}^{2}(\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})\rightarrow H^{2}(\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})$
Recall the pure Hodge structure $\partial H^{1}(\theta\phi)$ defined in (44). It is a direct summand of $H^{1}_{\mathop{\rm Eis}\nolimits,c}(\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})$, hence a quotient Hodge structure of the kernel of the second arrow in the above exact sequence. A reasoning analogous to the proof of Theorem 7.1, but replacing the use of Corollary 5.15 by the use of Corollary 5.16, and pullback by pushout via the epimorphism onto $\partial H^{1}(\theta\phi)$, shows the following.
Proposition 7.4.
Let $k$ and $\phi$ be as in Theorem 7.1.
  1. (1)
    For any level $K$ small enough, the short exact sequence (60) has a subquotient short exact sequence of Hodge structures
    (61) $0\rightarrow\partial H^{1}(\theta\phi)\rightarrow\tilde{E}_{0}^{\prime}\rightarrow H^{2}(\phi)_{!}\rightarrow 0$
  2. (2)
    Let $m^{\prime}(\phi,K)$ be the dimension of $I^{K}_{\theta\phi}$. Then the short exact sequence (61) induces an extension of $\overline{{\mathbb{Q}}}$-Hodge structures
    (62) $0\rightarrow\mathds{1}(1)^{\oplus m^{\prime}(\phi,K)}\rightarrow E_{0}^{\prime}\rightarrow H_{\phi}(-1)\rightarrow 0$
    where $H_{\phi}$ is the 1-dimensional $\overline{{\mathbb{Q}}}$-Hodge structure of type $(k,-(k+3))$ attached to $\phi$.
Remark 7.5.
Poincaré-Verdier duality provides a canonical isomorphism between boundary cohomology in degree 2 and the dual of boundary cohomology in degree 1. The above extension could have been obtained by applying such a duality to the extension of Theorem 7.1.
Corollary 7.6.
[…]
(63) $I^{K}_{\theta\phi}\rightarrow\mathop{\rm Ext}\nolimits^{1}_{\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}}}(H_{\phi}(-1),\mathds{1})$
7.2. Construction of biextensions
Consider the extensions $E_{\phi}(\Psi)$ constructed in the previous subsection. To verify the Bloch-Beilinson conjectures for our Hecke characters $\phi$, one has to show that there exists an element $\Psi$ of $I^{K}_{\phi}$ such that the extension $E_{\phi}(\Psi)$ is non trivial. This is the subject of work in progress by the authors, where verifying the desired non-triviality will rest crucially on the non-vanishing of a certain biextension height ([15], [9]) associated to $E_{\phi}(\Psi)$. Let us then define the input for such a height, i.e. a biextension in which $E_{\phi}(\Psi)$ sits.
We keep on looking at the local system $V_{k}$ on the Picard modular surface $\mathrm{S}_{K}$. As in subsection 4.3, we consider the Baily-Borel compactification $\mathrm{S}^{*}_{K}$ together with the open immersion $j$ of $\mathrm{S}_{K}$ and the closed immersion $i$ of the boundary $\partial\mathrm{S}_{K}$. To simplify notation, we will denote the latter by $\partial$.
We fix a decomposition of $\partial$ as the disjoint union of two sets of connected components
(64) $\partial=\Theta\bigsqcup\Sigma\,.$
For any $\triangle\in\{\Theta,\Sigma\}$, consider the pair of complementary, open resp. closed immersions
$j^{\triangle}:\mathrm{S}_{K}({\mathbb{C}})\hookrightarrow\mathrm{S}_{K}({\mathbb{C}})\bigsqcup\ \triangle\hookleftarrow\ \triangle:i^{\triangle}$
and the open immersion
$j_{\triangle}:\mathrm{S}_{K}({\mathbb{C}})\bigsqcup\ \triangle\hookrightarrow\mathrm{S}^{*}_{K}({\mathbb{C}})$
and define $i_{\triangle}:=j_{\triangle}\circ i^{\triangle}$. There are canonical exact triangles
$j_{\triangle,*}j^{\triangle}_{!}\rightarrow j_{*}\rightarrow i_{\triangle,*}i^{\triangle,*}j^{\triangle}_{*}\rightarrow$
For any local system $V_{\lambda}$ on $\mathrm{S}_{K}$, denote
$H^{i}_{c,\partial\setminus\triangle}(\mathrm{S}_{K}({\mathbb{C}}),V_{\lambda}):=H^{i}(\mathrm{S}_{K}^{*}({\mathbb{C}}),j_{\triangle,*}j^{\triangle}_{!}V_{\lambda}),$
$H^{i}_{\triangle}(\partial\mathrm{S}_{K}({\mathbb{C}}),V_{\lambda}):=H^{i}(\mathrm{S}^{*}_{K}({\mathbb{C}}),i_{\triangle,*}i^{\triangle,*}j^{\triangle}_{*}V_{\lambda}).$
We get a long exact sequence
(65) $\cdots\rightarrow H^{i}_{c,\partial\setminus\triangle}(\mathrm{S}_{K}({\mathbb{C}}),V_{\lambda})\rightarrow H^{i}(\mathrm{S}_{K}({\mathbb{C}}),V_{\lambda})\rightarrow H^{i}_{\triangle}(\partial\mathrm{S}_{K}({\mathbb{C}}),V_{\lambda})\rightarrow\cdots$
Moreover, there are canonical morphisms
(66) $i_{*}i^{*}j_{*}\xrightarrow{r_{\triangle}}i_{\triangle,*}i^{\triangle,*}j^{\triangle}_{*}$
such that the diagram
${j_{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${i_{*}i^{*}j_{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r_{\triangle}}$${i_{\triangle,*}i^{\triangle,*}j^{\triangle}_{*}}$
commutes.
Definition 7.7.
We define the subspaces
$I^{K}_{\phi,\Theta}\leq I_{\phi}^{K},\ I^{K}_{\theta\phi,\Sigma}\leq I_{\theta\phi}^{K}$
as the ones whose elements are supported on $\Theta$, resp. on $\Sigma$.
They acquire Hodge structures
$\partial H_{\Theta}^{2}(\phi),\ \mbox{resp.}\ \partial H_{\Sigma}^{2}(\theta\phi)$
by restriction of those defined in (43), resp. (44).
Proposition 7.8.
Let $\phi$ be a Hecke character of $F$ as in Theorem 7.1. Fix a decomposition
$\partial=\Theta\bigsqcup\Sigma\,.$
as in (64). There exist extensions (56) and (61)
(67) $0\rightarrow H^{2}(\phi)_{!}\rightarrow\tilde{E}\rightarrow\partial H_{\Theta}^{2}(\phi)\rightarrow 0$
(68) $0\rightarrow\partial H_{\Sigma}^{1}(\theta\phi)\rightarrow\tilde{E}^{\prime}\rightarrow H^{2}(\phi)_{!}\rightarrow 0$
and a canonical mixed $\overline{{\mathbb{Q}}}$-Hodge structure $\tilde{E}^{\sharp}$ fitting in a commutative diagram of mixed $\overline{{\mathbb{Q}}}$-Hodge structures, with exact rows and columns
[…]
Proof.
Consider the exact sequence
(69) $H^{1}_{\Sigma}(\partial\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})\rightarrow H^{2}_{c,\Theta}(\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})\rightarrow H^{2}(\mathrm{S}_{K}({\mathbb{C}}),V_{k})\rightarrow H^{2}_{\Sigma}(\partial\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})$
obtained from the extension of scalars of the long exact sequence (65). We have inclusions
$I^{K}_{\phi,\Theta}\hookrightarrow H^{2}_{\Theta}(\partial\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}}),\ \mbox{resp.}\ I^{K}_{\theta\phi,\Sigma}\hookrightarrow H^{1}_{\Sigma}(\partial\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})$
which factor as the composition of
$I_{\phi,\Theta}^{K}\hookrightarrow H^{2}(\partial\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}}),\ \mbox{resp.}\ I_{\theta\phi,\Sigma}^{K}\hookrightarrow H^{1}(\partial\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})$
and of the morphisms induced by (66). By semi-simplicity of $H^{1}_{\Sigma}(\partial S_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})$), there exists a projection
$H^{1}_{\Sigma}(\partial\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})\twoheadrightarrow I^{K}_{\theta\phi,\Sigma}$
Taking pushout along the latter, the sequence (69) yields an exact sequence
$0\rightarrow\partial H_{\Sigma}^{1}(\theta\phi)\rightarrow\tilde{\tilde{E}}_{\Sigma}\rightarrow H^{2}(\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})$
Denoting by $\tilde{\tilde{E}}$ the pullback of
$H^{2}(\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})\rightarrow H^{2}(\partial\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})$
along
$I^{K}_{\phi,\Theta}\hookrightarrow H^{2}(\partial\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})$
we obtain, by pullback via $\tilde{\tilde{E}}\rightarrow H^{2}(\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})$, an extension
(70) $0\rightarrow\partial H_{\Sigma}^{1}(\theta\phi)\rightarrow\tilde{E}_{\Sigma}\rightarrow\tilde{\tilde{E}}\rightarrow 0$
An analogous pullback applied to the extension (56) yields an extension $\tilde{E}$ as in in (67). Pullback of (70) along the induced morphism
$\tilde{E}\rightarrow\tilde{\tilde{E}}$
provides an extension
$0\rightarrow\partial H_{\Sigma}^{1}(\theta\phi)\rightarrow\tilde{E}^{\sharp}\rightarrow\tilde{E}\rightarrow 0$
which gives the right four-term column of the diagram in the statement. ∎
Corollary 7.9.
Let $\phi$ be a Hecke character of $F$ as in Theorem 7.1. Fix a decomposition
$\partial=\Theta\bigsqcup\Sigma\,.$
as in (64) and let $m(\phi,K,\Theta)$ be the dimension of $I_{\phi,\Theta}^{K}$ and $m^{\prime}(\phi,K,\Sigma)$ the dimension of $I_{\theta\phi,\Sigma}^{K}$. There exists a canonical diagram of mixed $\overline{{\mathbb{Q}}}$-Hodge structures
[…]
Proof.
By Corollary 5.15, the pure $\overline{{\mathbb{Q}}}$-Hodge structure $\partial H_{\Theta}^{2}(\phi)$ is a direct sum of copies of the 1-dimensional $\overline{{\mathbb{Q}}}$-Hodge structure $I_{k,-k-3,\overline{{\mathbb{Q}}}}$. By tensoring the diagram of Proposition 7.8 by the dual of $I_{k,-k-3,\overline{{\mathbb{Q}}}}$, we obtain the statement. ∎
To conclude our construction, we need to recall the concept of biextension (cfr. for example [9], Def. 2.1) in our setting.
Definition 7.10.
Let $H$ be a pure $\overline{{\mathbb{Q}}}$-Hodge structure of weight $-1$. A biextension of $H$ is a $\overline{{\mathbb{Q}}}$-mixed Hodge structure $B$ together with an identification of its associated weight-graded semisimple Hodge structure
$\mathop{\rm Gr}\nolimits^{W}B\simeq\mathds{1}(1)\oplus H\oplus\mathds{1}$
By a similar reasoning to the one leading to Corollaries 7.2 and 7.6, we obtain the following.
Corollary 7.11.
Let $\phi$ be a Hecke character of $F$ as in Theorem 7.1. For any decomposition
$\partial=\Theta\bigsqcup\Sigma\,.$
as in (64) and any
$(\Psi,\Psi^{\prime})\in I_{\phi,\Theta}^{K}\times I_{\theta\phi,\Sigma}^{K}$
the diagram of Corollary 7.9 induces a biextension
(71) $B_{\phi}(\Psi,\Psi^{\prime})$
of $H_{\phi}(-1)$.
Acknowledgements
The authors would like to thank for their hospitality and wonderful working environment, the Georg-August Universität Göttingen, Technische Universität Dresden, Mathematisches Forschungsinstitut Oberwolfach (MFO), Max Planck Institute für Mathematics, Bonn, Germany, the Institut de Recherche Mathématique Avancée (IRMA), Strasbourg, the Laboratoire de Mathématiques d’Orsay (LMO) and the Institut de Mathématiques de Bourgogne (IMB), Dijon where much of the discussion and work on this project was accomplished.
[…]